A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.12

Title word cross-reference

(L^{p}, L^{p}) [95]. (SO_{p+q}, SO_{p} × SO_{q}) [201]. (Sp_{2n}, O(V)) [477]. ∗ [157, 360]. 1 [111]. 2 [18, 149, 532]. 25 [170]. 3 [429, 105]. 4 [170]. 5 [197, 178, 492]. a [522]. A(G) [501]. A_{p}(G) [222]. A_{θ} [248]. ℵ_{1} [187]. ax^{p} + by^{p} = cz^{2} [324, 324]. b [247]. (M_{2}(K), t) [157]. (Sp(2, R), SL(2, C)) [135]. modp [312]. C^{∗} [207, 59, 321, 260, 63, 280, 380, 70, 101, 94, 304, 80, 363, 360, 109]. C^{1} [148]. CAR [129]. CAT(0) [282]. cb [406]. CM [100]. E [187]. F_{1} [41]. floor n^{c} [73, 73]. G [44, 364]. G/B [458]. G_{2} [303]. GL(2) [12]. GL(3) [193]. GL(n, F) [362]. GL_{n} [358, 55]. GSp(2) × GL(2) [502]. H [414, 354, 163]. H^{2} [203, 204]. H^{∞} [466]. H^{p} [415]. K [7, 248, 390, 88, 81, 281]. k(n) [414]. K = 0 [160]. K_{0} [50]. K_{1} [101]. L [128, 394, 285, 287, 100, 502, 401]. L^{1}(ω) [353]. L^{2} [31, 275, 405]. l^{∞} [437]. l_{1}^{m} [43]. L_{0} [409]. L_{p} [251]. λ^{[O(n^{15/14+o(1)})]} [498]. F_{8} [170]. P^{1} × P^{1} [239]. P^{2} [178]. P^{7} [531]. Q [257, 296]. R^{5} [316]. S^{3} [520]. Z [494, 550]. Z[\sqrt{14}] [209]. P^{2} [10]. Q [149]. CR [307, 385]. Z [453]. ⇕ [370]. √ [171]. R^{n}

[496]. μ [484, 361, 296]. n [78, 467, 479]. P [389, 24, 337, 199, 112, 417, 426, 51, 27, 469, 428, 391, 393, 290, 492, 503, 174]. P (n) [414]. ¶^{n} [92]. p_{g} = q = 2 [182]. ϕ_{1}x_{1}^{2} + ⋯ + ϕ_{n}x_{n}^{2} = 0 [29]. q [130, 23]. R [71]. R^{d} [230]. ρ = 1 [531]. S = 0 [160]. s = 1 [100]. SL(2) [290]. SL(3, C) [229]. SL_{3}(Z)

[517]. SL_{3}(Z) [153]. SL_{n} [384, 356]. Sp(2, R) [133]. Sp_{8} [214]. Spin(7) [303]. SU(2) [17]. SU(2, 1) [547]. SU_{n} [332]. tA_{2} [88]. τ [475]. U(2, 1)(F_{0}) [456]. U_{3} × GL_{2} [546]. W [37]. W^{∗} [244].

-Adic [51, 469, 492, 171, 503, 174, 24, 199, 112, 426, 27, 290]. -adique [199, 112]. -Algebra [94, 260]. -Algebras [207, 59, 321, 360, 63, 280, 380, 70, 101, 304, 109, 363, 80]. -Analogs [23]. -ary [247]. -Automorphisms [360]. -Bloch [484]. -Cohomology [414]. -Concave [391]. -Configurations [281]. -Convex [391]. -Curves [149]. -Diagonal [71]. -Dimensional [467, 479]. -Epic [385, 307]. -Estimate [405]. -Extensions [393]. -Fields [100]. -Functions [128, 394, 287, 100, 502, 285]. -Graded [364]. -Groups [50, 37]. -Homology [257, 248]. -Ideals [389]. -Integral [130]. -Invariant [522]. -Invariants [296]. -Laplacian [337]. -Manifolds [197, 429]. -Maximal [415]. -modules [550, 494]. -Multiplier [406]. -Orthogonal [130]. -Primary [370]. -Punctured [401]. -Rank [532]. -Representations [17]. -Rings [187]. -Spaces [414, 251]. -Sphere [105]. -Stable [453]. -Subvarieties [157]. -Summing [391]. -Theory [7, 88, 81, 390]. -Torsion-Free [414]. -TRO [244]. -Tuples [78]. -Typical [428]. -Valued [437]. -Varieties [44]. -Yang [417].

2-Generator [547].

3-Folds [531]. 3-Manifolds [188].

Abelian [192, 448, 83, 205, 312, 223]. Absolute [307, 385]. Acting [88]. Actions [447, 151, 205, 231, 65, 226, 262, 235, 302, 465]. Acyclic [439]. Adic [185, 51, 469, 492, 171, 503, 174, 24, 199, 112, 426, 27, 290]. adique [199, 112]. Adjacency [265, 468]. adjoint [294]. adjointe [294]. Admissibility [404]. Admissible [203, 204]. Admitting [429]. AF [464, 63]. AFD [244]. Affine [327, 427, 344, 124, 177]. Alexander [188]. Algebra [464, 406, 129, 222, 501, 94, 460, 261, 527, 528, 529, 144, 317, 260]. Algebraic [266, 494, 550, 338, 512, 210, 41, 124, 234, 44, 410]. Algebras [207, 327, 366, 495, 31, 539, 59, 426, 321, 26, 113, 229, 248, 360, 51, 52, 349, 205, 188, 260, 11, 213, 63, 280, 380, 70, 393, 101, 94, 124, 459, 373, 304, 80, 109, 326, 319, 171, 363, 191, 453, 364, 530, 112]. Algèbres [395, 530, 112]. algébrique [410]. Algorithm [10]. Algorithms [168]. Aligned [363]. Almost [187, 536]. Almost-Free [187]. along [158, 424]. Alternating [369]. Amalgamation [71]. Amenability [266, 406, 222]. Amibes [432]. Ampère [33, 46]. Ampère-type [33]. Ample [486]. Amplitude [242]. Analogs [23]. Analysis [375, 76, 476]. Analytic [138, 495, 158, 424, 219, 341, 196, 389, 517]. annihilator [220]. Annihilators [342]. Annular [429]. Apolar [338]. Appell [41]. Application [345]. Applications [138, 339, 497, 458, 491, 502, 54, 362]. Approach [446, 8, 410, 382]. approche [410, 382]. Approximants [328]. Approximate [430]. Approximating [282]. Approximation [138, 313, 489, 241, 374]. Approximations [121, 543]. Archimedean [502]. Area [423]. Arising [149, 288, 11, 70, 81]. arithmetic [137]. arithmétiques [137]. Arrangement [527]. Arrangements [189]. ary [247]. Ascent [485]. ASH [453]. Associated [464, 416, 159, 522, 227, 150, 249, 524, 61, 391, 190, 180, 169, 319, 153]. associées [395]. Asymptotic [437, 310, 88, 181]. Asymptotically [39, 292, 405]. Asymptotics [230]. asymptotique [480]. Auslander [485]. auteurs [56, 103, 155, 206, 264, 318, 365, 422, 481, 548]. Author [56, 103, 155, 206, 264, 318, 365, 422, 481, 548]. Automorphic [32, 61, 179]. Automorphism [81]. Automorphismes [177]. Automorphisms [360, 70, 177]. Avoiding [83, 211]. Axiom [284].

Backtracking [400]. ball [331]. Banach [291, 299, 368, 399, 351]. Base [178, 12, 456]. Bases [418]. Baskakov [137]. Behavior [310, 142, 242, 471, 98]. Bending [270]. Bergman [147, 219, 490]. Bernstein [368]. Berry [250]. Besov [152]. Best [227, 241]. between [273]. Bi [347]. Bi-Invariant [347]. Bifurcations [141]. bigèbre [395]. Billiard [212]. Biprojectivity [144]. Birch [278]. Birkhoff [253]. Bivariate [77, 121]. bivariés [121]. Bloch [484, 55]. Blowing [425]. Blowing-Down [425]. Blowing-Up [425]. Bodies [423, 343, 362]. Bol [377]. bornées [93]. Bottle [334]. boule [331]. Bound [20]. Boundary [15, 146, 429, 232, 233, 46]. Bounded [391, 490, 46, 93]. Bounds [448, 85, 541, 100, 491, 3]. Brackets [191]. Braid [79]. Braids [189]. Branched [17]. Branching [450, 290]. Branges [507]. Brownian [8, 41, 273]. Building [72]. Buildings [48, 521, 88]. Bump [238]. Bundle [329]. Bundles [356, 53, 190, 102].

Cabled [383]. Calabi [531]. calculations [442]. calculs [442]. Calculus [8, 466]. Calderón [416]. Canonical [49, 418, 326]. Cantor [110, 308, 70]. Cantor-Type [110]. Carathéodory [152]. Cardinal [389]. Cardinality [187]. Carleson [325]. Carnot [152]. Cartan [419, 195]. cas [237, 112, 331]. Case [426, 456, 237, 112, 331]. Cases [421]. Casselman [358]. CAT [521]. Categorical [139]. Categories [485, 14]. Categorification [474]. Category [125]. Cebysev [496]. Cell [444]. Central [260, 216]. Centralizers [345]. Certain [463, 176, 61, 279, 70, 372, 101, 214, 33]. certaines [33]. Chain [439]. Chains [511]. Chandra [361, 382]. Change [12, 456]. changement [456]. Changes [186]. Character [150]. Characteristic [413, 428]. Characterization [515, 458]. Characterizing [487]. Characters [18, 35, 545, 30]. Chen [407]. Chern [30]. Chief [131]. Choquet [399]. Circle [231]. Circles [538]. Circular [508]. Class [377, 404, 369, 342, 100, 515, 443, 544, 246]. Classes [426, 70, 372, 517]. Classical [433, 537, 67, 523, 27, 242, 372, 503, 518, 47, 530, 253]. Classification [199, 425, 367, 240, 63, 470, 101, 109, 364, 199]. classiques [530]. Close [538, 362]. Closed [138, 495, 398, 452, 295]. Closure [201]. Closures [16]. CM [526]. CM-Connectivity [526]. Co [266, 354, 163]. Co- [354, 163]. Co-Amenability [266]. Coalescing [273]. Coassociative [354, 163]. Cochain [154]. Cocycles [190]. Codes [65]. Coefficient [186]. Coefficients [522, 285, 499, 524, 306]. Cogrowth [400]. Cohen [440]. Cohomology [346, 414, 275, 310, 356, 333, 407, 190, 458, 231, 459, 517]. Coinvariant [229]. Coinvariants [435]. Colored [474]. Combinatorics [143]. Commutant [172]. Commutative [440, 81]. Commutator [219]. Commuting [78, 249, 302]. Compact [207, 347, 269, 259, 42, 43, 134, 304, 261, 442]. compactes [442]. Compactification [472]. Compactness [113]. Comparison [7]. Comparisons [108]. Compatible [312]. Complements [17]. Complete [5, 487, 360, 95, 54, 298, 137, 137]. Complex [183, 217, 200, 201, 114, 511, 151, 190, 215, 153, 46, 154]. Complexes [14, 282]. Components [482, 115]. comportement [98]. Composition [484]. Computational [528]. Computations [388]. Concave [391]. Concentration [510]. Conclusion [313]. Conditions [15, 146, 311, 466, 13]. Conductor [54]. Cone [301, 486, 522, 398, 295]. Cone-Monotone [301]. Cones [370]. Configurations [281]. Conformal [510]. Congruence [517, 504]. Conjecture [508, 359, 483, 461, 309, 421, 514, 278, 224, 352, 336]. Conjugacy [426, 430]. Connected [482, 115, 344]. Connectedness [259]. Connections [327, 329, 259, 28]. Connectivity [526]. Constant [351, 443]. Constants [227, 164]. Constructing [320, 431]. Construction [168, 517]. Contact [317]. Contained [383]. Context [452]. Continua [513]. Continuation [412]. Continued [153]. Continuity [301]. Continuous [366, 413, 59, 94, 154]. Contractions [364]. Convergence [369, 113, 328]. Convex [76, 423, 511, 313, 391, 343]. Convexité [137]. Convexity [137, 217, 159, 195]. Convolution [311, 113]. Coordinates [444]. Corepresentations [266]. corps [362]. Correction [85]. Correlation [174]. Correlations [216]. Correspondence [102]. Correspondences [31]. Corrigendum [550, 9, 74]. Cosingular [314]. Countable [513, 519]. Counting [85, 3]. courbes [184, 237]. Covering [284, 250, 42, 535]. Coverings [284, 108]. Covers [17, 320, 428]. Cowen [240]. Coxeter [521]. CR [473]. Critical [276]. Cross [344, 64, 547]. Cross-Ratios [547]. Cross-Section [64]. Crossed [205]. Cubes [277, 117]. Cubic [483, 142, 69, 12, 153]. Cubical [408]. Cuntz [129, 205]. Current [459]. Curvature [472, 292, 443]. Curve [265, 329, 170]. Curved [275]. Curves [75, 454, 158, 424, 445, 149, 350, 61, 279, 381, 230, 234, 215, 296, 184, 237]. Cuspidal [412, 346]. Cycle [255]. Cycle-Level [255]. Cycles [141, 533, 181]. Cyclic [387, 17, 418, 342]. Cyclotomic [183].

dans [118, 237, 331]. Darboux [68, 317]. Data [412]. Decay [311, 202, 216]. Decomposability [349]. Decomposition [419, 195, 244]. Decompositions [92, 28, 354, 163]. Decreasing [447]. Dedekind [100, 4, 74]. Défaut [237]. Defectivity [463]. Defined [25, 129]. Definite [335, 241]. Deformations [303]. Degeneration [510]. Degree [149, 342, 491, 393]. Dehn [252]. Dense [308]. Density [510, 352]. Deny [399]. Dependent [15, 146]. deployed [355]. déployés [355]. Depth [18]. Depth-Zero [18]. Derivations [191]. Derivatives [238]. Derived [133]. Descending [75]. Descent [508]. Describing [403]. Determined [447]. Deviation [77]. d’exponentielles [432]. d’homotopie [395]. Diagonal [71]. Diagonals [123]. Diagrams [265]. Diametrically [351]. Dieudonné [419]. Difference [288, 505]. Differences [343]. Differentiability [301, 243]. Differential [146, 84, 166, 221, 295]. Diffusions [509]. Digit [247]. Dilations [78]. Dilogarithm [119]. Dimension [330, 212, 489, 379, 451, 152]. Dimensional [347, 228, 467, 211, 72, 326, 479, 535, 181]. Dimensions [85, 450, 44, 95, 3]. Ding [367]. Diophantine [208, 446, 489]. Dirichlet [457, 352]. Disconnected [134]. Discrete [288, 230, 171, 545, 54, 246, 47]. discriminant [19]. Distance [387, 429, 392]. Distribution [394, 506, 348, 549]. Distributions [355, 35, 319]. Distributive [504]. Divisible [169]. Division [256, 393, 256]. Domains [145, 146, 9, 315, 236, 96, 173, 402, 1]. Donaldson [102]. d’ordre [361]. Double [183, 501]. Douglas [240]. Down [425]. Dual [501, 477, 353]. Duality [271, 273]. d’un [199]. d’une [331, 534]. Dunford [34]. Dupin [316]. Duplication [41]. Dyer [278]. Dynamical [321].

Effective [419, 151]. Eigenfunction [146]. Eigenparameter [15, 146]. Eigenvalue [334]. Eigenvalues [390]. Einstein [537].

Eisenstein [412, 225]. Elementary [111, 224]. Elements [501, 71, 245, 545, 218, 218]. Elliptic [75, 145, 138, 445, 388, 398, 332, 350, 279, 215, 296, 237]. elliptiques [237]. Embeddings [307, 122, 254]. Endomorphism [379, 373]. Endomorphisms [129, 460]. Ends [9, 1]. Energy [230, 242]. Enriques [265]. Epic [385, 307]. epsilon [98]. epsilon-pair [98]. Equation [420, 142, 399, 319]. Equations [337, 145, 91, 208, 15, 146, 138, 446, 6, 107, 236, 471, 127, 33, 324, 33, 324]. Equivalence [367]. Equivariant [57, 139, 122, 262, 184]. équivariante [184]. Erdos [487, 223]. Ergodic [227]. Erratum [434, 549, 166, 462]. espace [177]. espaces [118]. Esseen [250]. Essential [44]. Estimate [405, 21]. Estimates [25, 292, 212, 335, 451, 38, 173]. Eta [4, 74]. Euclidean [475, 39, 209, 210, 319]. Euler [183, 276, 41, 371]. Evaluations [41]. Eventual [196]. Every [308]. Exact [452]. Exactly [156, 434]. Examples [531]. Exceptional [32, 421, 277, 224, 411, 117]. Excluding [247]. exemple [456]. Existence [337]. Expansion [283]. Expansions [146, 475, 416]. Explicit [100, 171]. Exponential [91, 498, 279, 216, 93]. exponentiels [93]. Exponents [489, 374]. Ext [70]. Extended [327]. Extension [473, 339]. Extensions [413, 445, 139, 418, 398, 280, 37, 393, 460, 295, 54]. Extremal [334, 179].

Face [526, 527]. facteurs [98]. Factor [131]. Factorization [299, 172]. Factorizations [436, 86, 533]. Factors [58, 113, 98]. failure [237]. Families [446]. Family [512]. Farey [464]. Fast [204]. Fat [239, 10]. Features [159]. Feichtinger [359]. Fekete [397]. Fiber [106, 522, 370]. Fibre [189]. Fibre-Type [189]. Field [59, 20, 342, 290, 492, 460, 216]. Fields [75, 165, 475, 445, 417, 437, 388, 532, 350, 100, 37, 373, 225, 54, 153, 174]. Fillings [429]. Filters [438]. Finding [247]. Finer [367]. Finite [165, 320, 78, 229, 150, 440, 379, 72, 191, 519]. Finite-Dimensional [72]. Finitely [131, 363, 444]. Finsler [160, 544]. First [513, 334]. Fixed [242]. Flag [16, 443]. Flat [292, 259, 28, 443]. Flats [521, 282]. Floer [17]. Flow [493, 274]. Flows [270, 390, 253]. Focus [221]. Folds [531]. Foliations [136]. fonction [361]. fonctions [137]. Forcing [513]. Form [146, 253, 73]. Formality [262]. forme [73]. Formed [58]. Forms [283, 208, 338, 523, 61, 457, 515, 306]. Formula [35, 62, 358, 184]. Formulae [247, 285, 41]. formule [184]. fortes [121]. Four [330]. Fourier [433, 283, 406, 328, 501, 261, 30, 81, 306, 144]. Fractals [444]. Fractional [506]. Fractions [153]. Frame [168]. Framed [53, 116]. Frames [359, 53, 478, 126]. Fredholm [274, 38]. Free [414, 66, 187, 205, 304]. Free-Product [304]. Freeness [71]. French [118, 137, 530, 199, 456, 218, 184, 98, 237, 121, 112, 256, 355, 177, 331, 33, 361, 362, 324, 93, 442, 19, 410, 73, 294, 382]. Frey [446]. Friction [403]. Fubini [251]. Full [420]. Fully [145]. Function [175, 186, 378, 219, 85, 92, 203, 204, 41, 371, 500, 233, 373, 460, 225, 543, 529, 4, 74, 3, 361]. Functional [466, 319]. Functions [5, 110, 128, 394, 413, 301, 495, 348, 284, 438, 258, 341, 238, 202, 286, 239, 196, 328, 287, 335, 457, 42, 100, 36, 381, 133, 135, 502, 94, 352, 243, 169, 411, 535, 452, 536, 89, 505, 46, 90, 137, 549, 285]. Functor [133]. Functorial [354]. Functors [40, 439]. Fundamental [510, 401]. Furstenberg [430].

Gabor [359]. Galois [208, 7, 149, 467, 312, 393, 431]. gamma [137]. Gap [417]. Gaps [283]. Gaskets [535, 536]. Gauge [363]. Gauge-Invariant [363]. Gauss [69]. Gaussian [437, 173]. Gelfond [305]. General [145, 514, 108, 459, 492, 293, 503, 364]. Generalizations [223]. Generalized [327, 346, 167, 512, 99, 516, 289, 281, 254, 172]. Generated [507, 157, 131, 401]. Generating [203, 204]. Generation [366]. Generator [547]. Generic [418, 332]. Genus [454, 381, 170, 462, 2]. Geodesic [241]. Geodesics [232]. Geometric [521, 343, 22]. Geometry [520, 398, 28, 409, 428, 221, 441, 544, 105]. Ginzburg [452]. Girsanov [509]. Given [335]. Global [508, 445, 197, 379]. Goldbach [277]. Gordon [382]. Graded [139, 522, 364]. Graph [175, 539, 45]. Graphical [211]. Graphs [375, 425, 250, 132, 134, 400, 363]. Grassmannian [162]. Grassmannians [120]. Green’s [233]. Gross [525]. Grothendieck [57]. Ground [403]. Group [435, 436, 139, 18, 340, 269, 420, 150, 342, 151, 288, 64, 232, 233, 304, 65, 226, 401, 465, 199]. groupe [199]. groupes [218, 355, 93]. Groupoids [261]. Groups [24, 537, 128, 82, 266, 57, 347, 521, 139, 387, 320, 311, 50, 32, 194, 67, 323, 523, 350, 27, 271, 469, 205, 189, 514, 43, 108, 37, 344, 134, 372, 124, 492, 44, 88, 215, 262, 171, 503, 246, 22, 126, 55, 518, 47, 355, 93, 478]. Growth [378, 42].

Haar [323]. Hamiltonian [403, 191]. Hankel [490]. Hardy [376, 415, 21, 172]. Harish [361, 382]. Harish-Chandra [361, 382]. Harmonic [375, 300, 200, 444]. Hausdorff [300, 341, 212, 152]. Hauteur [480]. Heat [311, 233, 143]. Hecke [394]. Heegner [445, 480]. Heisenberg [420, 288, 315, 232, 233]. Hermitian [499, 468, 298]. heuristic [382]. heuristique [382]. Hexagons [254]. Hierarchy [68]. Higgs [298]. High [211].

High-Dimensional [211]. Higher [473, 386, 158, 424, 39, 349, 277, 535, 181, 363, 55]. Higher-Rank [363]. Hilbert [522, 92, 532, 239, 465]. Hitchin [102]. Hochschild [154]. Hodge [263]. Hölder [472, 413, 238]. Holomorphic [366, 31, 42]. Homeomorphisms [340, 519]. Homogeneous [537, 25, 330, 178]. Homology [17, 257, 248, 408, 188, 198, 55]. Homotopy [163]. Hopf [188, 180]. Horn [9, 1]. Horn-Like [9, 1]. Hyperbolic [76, 378, 197, 429, 304, 169, 253, 256]. hyperbolique [256]. Hypercontractivity [286]. Hypergeometric [192, 41]. Hyperinvariant [455]. Hyperoctahedral [436]. Hyperplane [189, 527, 162]. Hyperspaces [496]. Hypersurfaces [316]. Hypoelliptic [347].

Ideal [299, 205, 80]. Ideals

[495, 389, 370, 353, 363, 402]. Ideas [408]. Idempotents [392]. II [424, 67, 342, 204, 226]. III [69]. IIIa [450]. Image [431]. imaginaire [331]. Imaginary [475, 331]. Immersions [515]. Implementers [129]. Imprimitivity [302]. Increments [437]. Indefinite [64]. Independence [451]. Index [56, 103, 155, 206, 264, 318, 365, 422, 481, 548]. Individual [247]. Induced [82, 271, 220]. Inductive [28]. inégalités [137]. Inequalities [376, 227, 397, 420]. Inequality [498, 137]. Inequivalent [86]. Inert [20]. Infinite [347, 497, 228, 260, 380, 500, 326, 536, 45]. Infinite-Dimensional [228]. Infinitely [169]. Infinitesimal [529]. Infinity [472, 141]. Injective [497]. Injectivity [148]. Inner [523, 203, 204]. Instabilities [417]. Integers [394, 494, 550, 475, 185, 210]. Integrability [253]. Integrable [27, 514, 180]. Integral [433, 130, 62, 546]. Intégrales [112, 362, 530]. Integrals [493, 25, 51, 41, 21, 325, 95, 530, 112, 362].

Interpolation [40]. Intersection [255, 504]. Intertwining [24, 287, 345, 518]. Invariant [537, 474, 347, 522, 427, 188, 35, 363, 89, 355]. invariantes [355]. Invariants [435, 7, 449, 26, 389, 87, 383, 529, 296]. Inverse [366, 451, 90]. Invertible [451]. Involutions [249, 401]. Involving [163]. Irrational [182]. Irrationalities [4, 74]. irreducible [93]. irréductibles [93]. Irreversible [321]. Isometric [78, 515, 465]. Isotropic [515]. Iterated [216]. Iterates [371]. Itself [340]. IV [455]. Iwasawa [508].

Jacobi [58]. Jacobian [410]. jacobiens [410]. Jacquet [82, 421]. Jet [396]. Jones [474]. Jordan [268, 319].

Kac [344, 64, 223]. Kähler [298]. Kampen [161]. Karamata [357]. Kato [54]. Kernel [532, 233]. Kernels [176, 311]. Kirillov [246]. Klein [334, 382]. Kloosterman [165]. Klyachko [492]. Knot [17, 383]. Knots [17, 383]. Knuth [245]. Kobayashi [102]. Kontsevich [89, 486]. Koras [57]. Kostant [23]. KP [68]. Kronecker [373, 460].

L [352]. L-Functions [352]. Lacunary [524]. Lagrangian [401]. Laguerre [512]. Lai [421]. Landau [452]. Landsberg [544]. Lang [309]. Laplacian [9, 337, 1, 420, 288, 315]. Laplacians [347]. Large [445, 397, 498, 142, 350, 429, 431]. Lattice [538]. Law [312, 216]. Laws [169]. Leading [186]. Least [77, 20]. Lebesgue [541]. Lefschetz [441]. Length [310]. Lens [87]. Lerch [543]. leurs [442]. Level [26, 349, 441, 352, 171, 255]. Levi [218]. Lie [537, 327, 395, 530, 112, 426, 269, 194, 51, 11, 93, 124, 459, 478, 245, 191, 22, 364]. Lifting [540, 467]. Lifts [477]. Like [9, 1]. Limanov [427]. Limit [141, 357, 216]. Line [276, 340]. Linear [15, 425, 178, 213, 108, 492, 471]. Lines [427]. Links [474]. Liouville [15, 186]. Lipschitz [236, 96, 173].

Littlewood [499, 488, 476]. Local [128, 448, 310, 540, 150, 357, 287, 288, 457, 317, 362]. Localization [14]. Locally [207, 134, 261, 191]. Locally-Finite [191]. locaux [362]. Log [257]. Logarithm [216]. Logarithmic [40]. Looped [354]. Loops [377, 163]. Lorentzian [292]. Low [438]. Low-Pass [438]. Lower [100]. Lusternik [125].

Macaulay [440]. Machin [247]. Machin-Type [247]. Mahler [119, 397]. Main [508]. Majorants [507, 203, 204]. Makar [427]. Manifold [232, 233]. Manifolds [104, 472, 473, 39, 275, 330, 292, 405, 151, 197, 188, 42, 429, 298, 317]. Map [340, 64]. Mapping [95]. Mappings [194, 114]. Maps [378, 148, 468, 451, 535, 125]. Marginal [72]. Marriage [541]. Martingales [226]. Matching [171]. Matrices [447, 268, 270, 341, 468, 123, 241]. Matrix [213]. Matroids [97]. Maximal [415, 227, 524, 393, 351, 456, 361]. maximaux [456]. Maximum [457, 36, 170]. Mazur [349]. Mean [423, 202]. Measurable [194]. Measure [217, 119, 397, 323]. Measures [46]. Merit [58]. Meromorphic [412, 138, 258]. Metaplectic [108]. Method [183, 305]. Methods [397]. Metric [413, 334, 304, 152]. Metrics [537, 160, 443, 544, 298]. Metrized [375]. Mikhlin [466]. Mikhlin-Type [466]. Mild [516]. Mills [298, 417]. Milnor [106]. Minimal [423, 483, 533, 390, 230, 70, 220]. Minimum [36]. Minkowski [343]. Mixed [263]. Mixing [516]. Model [156, 434, 403, 203, 204, 358]. Models [291, 308, 492, 542, 439, 452]. Moderate [177]. modérés [177]. Modular [283, 208, 6, 26, 306]. Module [7, 32, 514]. Modules [82, 494, 396, 66, 133, 373, 460, 38, 30, 550, 220]. Moduli [482, 486, 356, 259, 28]. Modulo [508, 498, 488]. Modulus [111, 36]. Moment [300, 130]. Monge [33, 46]. Monoidal [439]. Monotone [97, 301, 360]. monotonicity [137]. monotonie [137]. Moody [344, 64]. Mordell [350]. Morse [116]. Motions [273]. Motivic [323, 411]. Moving [168, 478]. Multi [446, 79]. Multi-Frey [446]. Multi-Parameter [446]. Multi-Sided [79]. Multiple [263]. Multiplication [454, 219, 499, 215]. Multiplications [314]. Multiplicative [202]. Multiplicity [337, 322, 178, 526]. Multiplier [406, 466, 222, 80]. Multipliers [147, 415]. Multisymplectic [235]. Multivariable [438]. Mutation [156, 434].

Near [472, 46]. Nearly [369]. Negative [472, 491]. Negatively [275]. Nest [52]. Nested [42]. Networks [211]. Neumann [9, 31, 1, 322, 33, 349, 236]. Newton [19]. Nilpotence [377]. Nilpotent [426, 201, 250, 133, 246, 294]. nilpotente [294]. Nilpotents [392]. Nombres [73]. Non [15, 509, 78, 330, 450, 440, 212, 491, 252, 451, 344, 223, 400, 319, 81, 237]. Non-Abelian [223]. Non-Backtracking [400]. Non-Commutative [440, 81]. Non-Commuting [78]. Non-Connected [344]. Non-Euclidean [319]. Non-Invertible [451]. Non-Linear [15]. Non-Negative [491]. Non-Orientable [252]. Non-reductive [330]. Non-Selfadjoint [450]. Non-Symmetric [509]. Non-Wandering [212]. Nonadjacent [475]. Noncommutative [251]. Noncommuting [435]. Nongraded [191]. Nonlinear [145, 322, 142, 471]. Nonmeager [308]. Nonnegative [268]. Nonsquare [164]. Nonstandard [353]. Norm [406, 313, 164]. Normal [253]. Normalized [287]. Normalizers [360]. Norms [200, 447]. Note [183, 461]. Nowhere [308]. Nuclear [391]. Number [475, 356, 532, 305, 374, 170]. Numbers [482, 300, 111, 100, 73]. Numerical [11, 213, 382]. numérique [382].

Octahedral [149]. Odd [426, 85, 3]. ODEs [549, 348]. One [270, 450, 352]. One-Level [352]. Onto [340]. Open [212]. Operads [439]. opérateur [382]. opérateurs [137]. Operator [493, 415, 140, 540, 219, 406, 325, 38, 144, 382]. Operators [433, 299, 376, 267, 274, 416, 200, 176, 269, 85, 398, 450, 524, 240, 287, 288, 391, 213, 43, 295, 172, 345, 518, 137, 484, 3]. Optimal [243]. Optimization [491]. Orbifold [102]. Orbit [16, 84, 166, 294]. Orbital [51, 530, 112, 362]. orbitales [530, 112, 362]. orbite [294]. Orbites [361]. Orbits [159, 201, 133, 465, 361]. Order [158, 424, 348, 39, 269, 221, 500, 519, 549, 361]. Ordered [50]. OrderExt [70]. Ordering [201, 189]. ordinaires [19]. ordinary [49, 106, 19]. Orientable [252]. Oriented [97]. Orlicz [164]. Orthogonal [348, 426, 26, 130, 271, 179, 549]. Orthogonality [384]. Oscillatory [95]. Osculating [386]. Other [13].

pôles [361]. packages [456]. packets [456]. Padé [328]. Painlevé [127]. Pair [201, 249, 135, 174, 98]. paires [98]. Pairings [333]. Pairs [477, 353]. Paley [476]. paquets [456]. Parabolic [329, 53, 102]. Parabolically [82]. Parallel [515]. Parameter [446]. Parametric [114]. part [331, 203, 204]. Partial [276, 360, 123]. Partie [331]. Partition [90]. Partitions [505]. Parts [506]. Pass [438]. Paths [97]. Patterns [83]. Pencil [182]. Perforated [50]. Periodic [348, 140, 535, 536, 282, 549]. Periodicity [539]. Periods [193]. Permutations [533, 134]. Perturbations [140, 450, 242]. Perverse [120]. Petits [534]. Petrie [297]. Pettis [34]. Phenomena [336]. Phi [371]. Phi-Function [371]. Pillai [91]. Pinched [472]. Planar [265, 9, 212, 1].

Plane [148, 343]. Planes [257]. Plurisubharmonic [46]. Poincaré [225]. Point [60, 450]. Points [75, 267, 445, 483, 538, 448, 92, 239, 10, 178, 441, 234, 170, 534, 480]. Pointwise [335]. Poisson [104, 541, 62, 191]. Polar [198]. Polarization [449]. Poles [258, 411, 518, 361]. Polydisk [490]. polyèdre [19]. Polygons [105]. Polyhedrality [228]. polyhedron [19]. Polylogarithms [263]. polynôme [256]. Polynomial [474, 313, 543, 256]. Polynomials [165, 76, 77, 58, 111, 348, 397, 499, 512, 524, 130, 179, 491, 516, 462, 383, 549, 2]. pondérées [112]. Positive [337, 394, 335, 391, 241]. Positivity [196]. Potential [22, 96]. Potentials [236]. pour [56, 103, 155, 206, 264, 318, 365, 422, 548, 530, 184, 121, 33, 382]. Power [342, 306]. Powers [311, 277, 505]. Pre [55]. Pre-Bloch [55]. Predual [222]. Preduals [391]. Prehomogeneity [518]. premiers [73]. Preserve [213]. Preserving [468]. Pressure [451]. Primary [370]. Prime [107, 20, 488, 342, 393, 305, 73]. Primes [498, 505]. Principal [329, 289, 290]. Principle [99]. Principles [457]. Pro [207, 124]. Pro- [207]. Pro-Affine [124]. Problem [346, 455, 403, 488, 277, 315, 226, 236, 33]. problème [33]. Problems [15, 186, 322, 90]. Process [72]. processes [121]. processus [121]. proches [362]. Product [66, 69, 21, 304, 325]. Products [276, 333, 205, 490, 401]. Projection [200, 219]. Projectionless [59, 109]. Projective [257, 407, 132, 234]. Projectively [443]. Prolongations [528]. Proof [224]. Proper [513, 235, 302]. Properties [417, 437, 485, 540, 219, 512, 197, 243, 95]. Property [284, 368, 349, 34, 441, 244, 504]. proto [395]. proto-bigèbre [395]. Pseudo [330, 478]. Pseudo-Riemannian [330]. Pseudoboundary [340]. Pseudoprimes [506]. Punctured [401]. Pure [189]. Purely [260, 380].

Quadratic [475, 107, 388, 141, 20, 221, 37, 373, 460, 352, 4, 74]. Quantum [266, 403, 458, 304]. Quasi [104, 49, 106, 341, 67, 205, 178, 518, 355, 19, 19]. quasi-deployed [355]. quasi-déployés [355]. Quasi-Free [205]. Quasi-Hausdorff [341]. Quasi-Homogeneous [178]. quasi-ordinaires [19]. Quasi-ordinary [49, 106, 19]. Quasi-Poisson [104]. Quasi-Split [67, 518]. Quasidiagonal [280]. Quasiregular [404]. Quaternionic [454, 32]. Quaternions [197]. Quelques [324, 442]. Quivers [418, 497]. Quotients [249, 43, 231, 410].

Radical [402]. Radius [213]. Radix [475]. Radix- [475]. Rallis [421, 23]. Ramanujan [48]. ramifié [237]. Ramified [54, 444]. Random [437, 279, 400, 216]. Range [213]. Ranges [11]. Rank [445, 18, 78, 539, 270, 532, 350, 492, 80, 363]. Rank- [18]. Ranks [94]. Rasmussen [474]. Rational [75, 185, 378, 36, 127]. Ratios [547]. Rayleigh [331]. Real [148, 388, 340, 40, 20, 80, 374]. Realization [521]. Realizations [66]. Reciprocity [312]. Recognition [226]. Recombination [156, 434]. Rectangular [244]. Reducibility [332, 289, 503]. réductifs [355]. Reduction [333, 235]. Reductive [327, 323, 355, 330]. Regular [254, 519, 218]. Regularity [415]. Regulated [373, 460]. réguliers [218]. Reidemeister [188]. Reiteration [40]. Related [506, 69, 325, 374, 191]. Relations [384, 245]. Relative [100, 317]. Remarks [476]. Renewed [357]. Renormalized [193]. Representation [299, 114, 546]. Representations [82, 208, 17, 18, 404, 320, 497, 149, 332, 467, 52, 53, 130, 27, 271, 312, 514, 491, 372, 477, 290, 431, 326, 401, 93, 199, 93, 199]. Residual [523, 214]. Residues [100, 306]. Resolution [49, 483, 44]. Resolutions [14]. Resolvent [267, 295]. Resonance [85, 3]. Resonances [140, 331, 331]. Respect [423]. Result [24]. résultats [324]. Results [275, 368, 322, 195, 357, 293, 503, 22, 324]. Reverse [286]. Richardson [499]. Riemann [184, 28]. Riemannian [275, 330, 33, 241]. riemannienne [33]. Riesz [217, 230]. Rigid [517]. Ring [522, 150, 458, 231, 526]. Rings [187, 210, 61, 379]. Roch [184]. Rook [367]. Rotation [248]. Rough [176]. Ruan [407]. Ruled [411]. Rules [290]. Russell [57, 6]. Russell-Type [6].

S [91]. S. [91, 461]. Same [258]. Satake [442]. Scaling [438]. Scattering [39, 242]. Scheme [154]. Schemes [338, 440, 297]. Schnirelman [305]. Schnirelmann [125]. Schrödinger [85, 140, 3, 142, 236, 471]. Schroeder [368]. Schubert [367]. Schwarzschild [405]. Searching [385]. Secant [386]. Second [348, 269, 459, 549]. Section [344, 64, 162]. Segal [68]. Segre [463]. Seiberg [87]. Selection [156, 434]. Self [211]. Self-Avoiding [211]. Selfadjoint [450]. Selmer [215]. Semi [433, 457, 242, 253, 237, 237]. Semi-Classical [433, 242, 253]. Semi-Dirichlet [457]. semi-stabilité [237]. Semi-stability [237]. Semigroup [13, 527]. Semilinear [315]. Semipositone [337]. Semiprojectivity [380]. Semisimple [135]. Sense [83]. Separable [368]. Separating [449]. Sequence [498, 260]. Sequences [300, 447, 526]. Series [412, 369, 289, 290, 225, 171, 545, 47]. Set [433, 308, 92, 212, 272]. Sets [5, 110, 138, 496, 277, 116, 441, 351, 117]. Setting [68]. Several [114, 511, 95].

Shalika [542, 358]. Sharing [258]. Sheaves [120]. Shells [423]. Shift [175, 493]. Shimura [115]. Short [165, 471]. Sided [79]. Sierpinski [535, 536]. Sieve [309, 397, 498]. Sign [186]. Similarity [240]. Simple [66, 320, 260, 11, 63, 280, 380, 101, 80, 109, 171, 220, 456]. simples [456]. Simultaneous [543]. Single [126]. singleton [456]. singletons [456]. Singular [337, 176, 25, 314, 21, 325, 226, 545, 236, 89, 317]. Singularités [19]. Singularities [265, 19]. Singularity [49, 106]. Size [272]. Sizes [283, 131]. Skein [188]. Slim [117]. Slow [203, 42]. Slowly [140]. Small [5, 107, 284, 29]. Smaller [75]. Smooth [284, 122, 440, 371]. Smoothness [249]. Snake [8]. Sobolev [413, 339]. Solutions [337, 138, 107, 29, 471, 127]. Solvability [288]. Solvable [234]. Solved [156, 434]. Some [537, 472, 91, 300, 495, 446, 159, 339, 275, 227, 311, 368, 421, 195, 41, 197, 293, 95, 152, 462, 503, 23, 2, 324, 442]. sommes [432]. sous [218]. sous-groupes [218]. Space [291, 486, 487, 399, 28, 490, 515, 21, 177]. Spaces [110, 537, 507, 307, 385, 299, 413, 147, 414, 415, 339, 84, 166, 540, 25, 368, 341, 259, 407, 251, 34, 42, 470, 351, 87, 304, 325, 354, 152, 172, 163, 164, 519, 465, 118, 484, 499]. Special [348, 549]. Spectra [398, 523]. Spectral [175, 493, 15, 267, 274, 219, 1, 9, 13]. Spectrum [67, 214, 45]. Sphere [520, 401, 105]. Spheres [143, 81]. Spherical [412, 135]. Spin [514]. Spinor [128, 405]. Split [128, 67, 523, 372, 518]. splitting [382, 382]. Spreading [291]. Square [267, 309, 27, 514, 374, 282]. Square-Integrable [27]. Squares [174]. stabilité [237]. stability [237]. Stable [541, 451, 545, 453]. Stably [109]. Stacks [482]. Stahl [461]. Standard [514]. Stark [336]. State [403]. States [129, 213]. Stationary [72]. Steinberg [344, 64, 479].

Steklov [376]. Stickelberger [336]. Stieltjes [328]. Stochastic [8, 72]. Stratification [60]. Stratified [116]. Strichartz [420]. Strictly [314]. Strong [121]. Strongly [132]. Structure [291, 7, 84, 166, 220, 429, 134, 80, 191, 444, 364, 465, 294]. Structures [268, 303, 205, 263]. Study [408, 343]. Sturm [15, 186]. Subalgebra [501, 504]. Subfactors [79]. Subgroups [16, 229, 517, 547, 218]. Subharmonic [397, 286, 457]. Sublinearity [13]. Submanifolds [241]. Submodules [360]. Subordination [511]. Subproportional [43]. SubRiemannian [520]. Subspace [455]. Subspaces [175, 507, 203, 204]. Subvarieties [157]. Sum [69, 123]. Summation [285, 62]. Summing [391]. Sums [165, 270, 277, 279, 117]. Supercuspidal [18]. Supersingular [381]. Support [335]. Supported [35]. Supports [271]. Surface [49, 106, 423, 483, 28, 180, 534]. Surfaces [510, 427, 259, 252, 293, 411, 102, 182]. Surgeries [252]. Swinnerton [278]. Swinnerton-Dyer [278]. Symbols [58, 479]. symétriques [118]. Symmetric [435, 16, 111, 509, 201, 499, 34, 135, 302, 118]. Symmetrizable [64]. Symplectic [18, 333, 28, 231, 105]. Systems [183, 321, 141, 312, 178, 221, 70, 180].

TAF [52]. Tail [541]. Tamagawa [482, 356]. Tangents [158, 424]. temperate [199]. Tempered [67, 477]. tempérées [199]. Tensor [66]. Ternary [208]. Tessellation [464]. Their [266, 386, 447, 239, 505, 523, 108, 442]. Theorem [387, 167, 419, 250, 251, 223, 44, 216, 162, 161, 23, 317, 302, 118]. Théorème [118]. Theorems [339, 525, 28]. Theoretic [22]. Theory [136, 348, 7, 227, 9, 60, 188, 116, 134, 502, 305, 88, 246, 255, 96, 81, 549, 1, 390]. Theta [32, 477, 225]. Thin [423]. Third [221]. Three [387]. Threefolds [57]. Time [142, 471]. Tits [521]. Toeplitz [172]. Tone [510]. Topologies [185]. Topology [408]. tordues [362]. Torelli [167]. Tori [327]. Toric [122, 255, 19]. toriques [19]. Toroidal [429]. Torsion [414, 66, 31, 448, 101, 98]. Torsors [482]. Torus [190, 262]. Totally [394, 268, 134]. TQFT [188]. Trace [35, 143]. Traces [547, 442]. Tracially [63]. Transfer [384, 530]. Transfert [530]. Transformations [300, 509, 68, 430]. transformées [442]. transforms [442]. Transitive [436, 86, 533]. Translates [359]. Transpositions [86]. transverse [294]. Trapping [242]. TRO [244]. Trotter [309]. Truncated [62]. Tsirelson [470]. Tuples [78]. Twisted [525, 345, 362]. twisting [98]. Two [377, 25, 450, 168, 489, 374]. Two-Weight [25]. Type [110, 48, 473, 247, 6, 25, 79, 466, 250, 440, 189, 470, 293, 244, 33]. Types [456, 469, 456]. Typical [428].

Ultraproducts [251]. Ultraspherical [416]. Unbounded [274, 315, 14]. unbranched [237]. Uniform [506, 217, 313]. Unipotent [218, 361]. unipotentes [361]. unipotents [218]. Uniqueness [542]. Unit [501, 171]. Unitarizability [503]. Unitary [159, 150, 151, 469, 372, 126]. Units [508, 388, 153]. Univalent [114]. Unramified [469]. Upper [3, 85, 20, 100].

Valuation [54, 402]. Valuative [136].

Value [348, 243, 549]. Valued [147, 415, 437, 94]. Values [202, 371, 100, 46, 4, 74]. Vanishing [28, 296]. Variables [435, 114, 511]. Variational [99]. Variations [41, 263]. variété [33]. Varieties [463, 192, 386, 158, 424, 16, 448, 367, 115, 122, 131, 44, 162, 255, 504, 89]. Variety [157]. Varying [140]. Vassiliev [383]. Vector [147, 356, 524]. Veronese [463, 386]. Very [431]. Via [457, 482, 208, 397, 188, 134, 526]. View [60]. Volume [423]. Voronoí [118].

Walks [211, 279, 400]. Wandering [212]. Waring [277]. Wave [420]. Wavefront [433]. Wavelet [359]. Wavelets [438, 126]. Weak [227, 380, 221, 441, 224, 171]. Weight [25]. Weighted [376, 339, 257, 425, 405, 60, 113, 53, 407, 112]. Weil [150, 350]. Well [54]. Weyl [82]. Which [308]. Whitney [110]. Whittaker [500, 133]. Whose [186, 447]. Width [423, 351]. Wild [272]. Wilson [68]. Winding [203, 204]. Without [508, 267]. Witt [419]. Witten [405, 87].

Yang [417, 298]. Yau [531].

zêta [137]. Zagier [525]. Zariski [161, 408, 162]. Zariski-van [161]. Zero [77, 18, 224, 80, 171]. Zeros [378, 258, 462, 2]. Zeta [378, 100, 381, 411, 319, 89, 137]. Zonotopes [97]. Zygmund [416].

References

Edward:1997:STN

[1] Julian Edward. Spectral theory for the Neumann Laplacian on planar domains with horn-like ends. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):232–262, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [9].

Stahl:1997:ZSG

[2] Saul Stahl. On the zeros of some genus polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 49(??):617–640, ???? 1997. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [462].

Froese:1998:UBR

[3] Richard Froese. Upper bounds for the resonance counting function of Schrödinger operators in odd dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 50(??):538–546, ???? 1998. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [85].

vanderPoorten:1999:VDE

[4] Alfred van der Poorten and Kenneth S. Williams. Values of the Dedekind eta function at quadratic irrationalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):176–224, February 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [74].

Aizenberg:2000:SCS

[5] Lev Aizenberg and Alekos Vidras. On small complete sets of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1): 3–30, February 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chan:2000:RTM

[6] Heng Huat Chan and Wen-Chin Liaw. On Russell-type modular equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (1):31–46, February 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chinburg:2000:CTG

[7] T. Chinburg, M. Kolster, and V. P. Snaith. Comparison of K-theory Galois module structure invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1): 47–91, February 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dhersin:2000:SCA

[8] Jean-Stéphane Dhersin and Laurent Serlet. A stochastic calculus approach for the Brownian snake. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1):92–118, February 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edward:2000:CST

[9] Julian Edward. Corrigendum to “Spectral Theory for the Neumann Laplacian on Planar Domains with Horn-Like Ends”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1):119–122, February 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1].

Harbourne:2000:AFP

[10] Brian Harbourne. An algorithm for fat points on P^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(1):123–140, February 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2000:NRA

[11] Chi-Kwong Li and Tin-Yau Tam. Numerical ranges arising from simple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):141–171, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mao:2000:CBC

[12] Zhengyu Mao and Stephen Rallis. Cubic base change for GL(2). Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):172–196, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjavi:2000:SOS

[13] Heydar Radjavi. Sublinearity and other spectral conditions on a semigroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):197–224, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tarrio:2000:LCC

[14] Leovigildo Alonso Tarrío, Ana Jeremías López, and María José Souto Salorio. Localization in categories of complexes and unbounded resolutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 225–247, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binding:2000:SPN

[15] Paul A. Binding, Patrick J. Browne, and Bruce A. Watson. Spectral problems for non-linear Sturm–Liouville equations with eigenparameter dependent boundary conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):248–264, ????

2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brion:2000:OCS

[16] Michel Brion and Aloysius G. Helminck. On orbit closures of symmetric subgroups in flag varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):265–292, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Collin:2000:FHK

[17] Olivier Collin. Floer homology for knots and SU(2)-representations for knot complements and cyclic branched covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):293–305, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cunningham:2000:CDZ

[18] Clifton Cunningham. Characters of depth-zero, supercuspidal representations of the rank-2 symplectic group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):306–347, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Perez:2000:SQO

[19] P. D. González Pérez. Singularités quasiordinaires toriques et polyèdre de Newton du discriminant. (French) [Quasiordinary toric singularities and Newton polyhedron of the discriminant]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 348–368, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Granville:2000:UBL

[20] Andrew Granville, R. A. Mollin, and H. C. Williams. An upper bound on the least inert prime in a real quadratic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):369–380, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miyachi:2000:HSE

[21] Akihiko Miyachi. Hardy space estimate for the product of singular integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):381–411, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:2000:GPT

[22] N. Th. Varopoulos. Geometric and potential theoretic results on Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):412–437, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallach:2000:SAT

[23] N. R. Wallach and J. Willenbring. On some q-analogs of a theorem of Kostant–Rallis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):438–448, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adler:2000:IRA

[24] Jeffrey D. Adler and Alan Roche. An intertwining result for p-adic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):449–467, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edmunds:2000:TWE

[25] D. E. Edmunds, V. Kokilashvili, and A. Meskhi. Two-weight estimates for singular integrals defined on spaces of homogeneous type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):468–502, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gannon:2000:LMI

[26] Terry Gannon. The level 2 and 3 modular invariants for the orthogonal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):503–538, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jantzen:2000:SIR

[27] Chris Jantzen. On square-integrable representations of classical p-adic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):539–581, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jeffrey:2000:SGM

[28] Lisa C. Jeffrey and Jonathan Weitsman. Symplectic geometry of the moduli space of flat connections on a Riemann surface: Inductive decompositions and vanishing theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):582–612, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ou:2000:SS

[29] Zhiming M. Ou and Kenneth S. Williams. Small solutions of ϕ_{1}x_{1}^{2} + ⋯ + ϕ_{n}x_{n}^{2} = 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):613–632, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walters:2000:CCF

[30] Samuel G. Walters. Chern characters of Fourier modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):633–694, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carey:2000:CNA

[31] A. Carey, M. Farber, and V. Mathai. Correspondences, von Neumann algebras and holomorphic L^{2} torsion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 695–736, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gan:2000:ATM

[32] Wee Teck Gan. An automorphic theta module for quaternionic exceptional groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):737–756, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanani:2000:PNP

[33] Abdellah Hanani. Le problème de Neumann pour certaines équations du type de Monge–Ampère sur une variété riemannienne. (French) [The Neumann problem for certain Monge–Ampèretype equations of Riemannian type]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):757–788, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaminska:2000:DPP

[34] Anna Kamińska and Mieczyslaw Mastylo. The Dunford–Pettis property for symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):789–803, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kottwitz:2000:DIT

[35] Robert E. Kottwitz and Jonathan D. Rogawski. The distributions in the invariant trace formula are supported on characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):804–814, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lubinsky:2000:MMM

[36] D. S. Lubinsky. On the maximum and minimum modulus of rational functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):815–832, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minac:2000:GUQ

[37] Ján Minác and Tara L. Smith. W - groups under quadratic extensions of fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):833–848, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sukochev:2000:OEF

[38] F. A. Sukochev. Operator estimates for Fredholm modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):849–896, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Christiansen:2000:HOS

[39] T. J. Christiansen and M. S. Joshi. Higher order scattering on asymptotically Euclidean manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 897–919, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Evans:2000:RIL

[40] W. D. Evans and B. Opic. Real Interpolation with logarithmic functors and reiteration. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):920–960, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:2000:AES

[41] Mourad E. H. Ismail and Jim Pitman. Algebraic evaluations of some Euler integrals, duplication formulae for Appell’s hypergeometric function F_{1}, and Brownian variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):961–981, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larusson:2000:HFS

[42] Finnur Lárusson. Holomorphic functions of slow growth on nested covering spaces of compact manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):982–998, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mankiewicz:2000:CGO

[43] Piotr Mankiewicz. Compact groups of operators on subproportional quotients of l_{1}^{m}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):999–1017, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reichstein:2000:EDA

[44] Zinovy Reichstein and Boris Youssin. Essential dimensions of algebraic groups and a resolution theorem for G-varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):1018–1056, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Urakawa:2000:SIG

[45] Hajime Urakawa. The spectrum of an infinite graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1057–1084, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xing:2000:CMA

[46] Yang Xing. Complex Monge–Ampère measures of plurisubharmonic functions with bounded values near the boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1085–1100, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2000:DSC

[47] Yuanli Zhang. Discrete series of classical groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1101–1120, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ballantine:2000:RTB

[48] Cristina M. Ballantine. Ramanujan type buildings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1121–1148, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ban:2000:CRQ

[49] Chunsheng Ban and Lee J. McEwan. Canonical resolution of a quasi-ordinary surface singularity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1149–1163, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:2000:POG

[50] George A. Elliott and Jesper Villadsen. Perforated ordered K_{0}-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 1164–1191, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herb:2000:OIA

[51] Rebecca A. Herb. Orbital integrals on p-adic Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1192–1220, ????

2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hopenwasser:2000:NRT

[52] Alan Hopenwasser, Justin R. Peters, and Stephen C. Power. Nest representations of TAF algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 1221–1234, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurtubise:2000:RWF

[53] J. C. Hurtubise and L. C. Jeffrey. Representations with weighted frames and framed parabolic bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 1235–1268, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spriano:2000:WRE

[54] Luca Spriano. Well ramified extensions of complete discrete valuation fields with applications to the Kato conductor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??):1269–1309, ???? 2000. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yagunov:2000:HHP

[55] Serge Yagunov. On the homology of GL_{n} and higher pre-Bloch groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52 (??):1310–1338, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2000:AII

[56] Anonymous. Author index — index des auteurs — for 2000 — pour 2000. Canadian Journal of Mathematics = Journal canadien de mathématiques, 52(??): 1339–1343, ???? 2000. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:2001:EGG

[57] J. P. Bell. The equivariant Grothendieck groups of the Russell–Koras threefolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):3–32, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2001:MFP

[58] Peter Borwein and Kwok-Kwong Stephen Choi. Merit factors of polynomials formed by Jacobi symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):33–50, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dean:2001:CFP

[59] Andrew Dean. A continuous field of projectionless C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):51–72, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fukui:2001:STW

[60] Toshizumi Fukui and Laurentiu Paunescu. Stratification theory from the weighted point of view. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):73–97, ????

2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khuri-Makdisi:2001:CAC

[61] Kamal Khuri-Makdisi. On the curves associated to certain rings of automorphic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):98–121, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levy:2001:TIP

[62] Jason Levy. A truncated integral of the Poisson summation formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 122–160, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:2001:CST

[63] Huaxin Lin. Classification of simple tracially AF C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):161–194, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mokler:2001:SMS

[64] Claus Mokler. On the Steinberg map and Steinberg cross-section for a symmetrizable indefinite Kac–Moody group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):195–211, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Puppe:2001:GAC

[65] V. Puppe. Group actions and codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):212–224, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Britten:2001:TPR

[66] D. J. Britten and F. W. Lemire. Tensor product realizations of simple torsion free modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):225–243, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldberg:2001:TSQ

[67] David Goldberg and Freydoon Shahidi. On the tempered spectrum of quasisplit classical groups II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):244–277, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helminck:2001:DTK

[68] G. F. Helminck and J. W. van de Leur. Darboux transformations for the KP hierarchy in the Segal–Wilson setting. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):278–309, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ito:2001:PRC

[69] Hiroshi Ito. On a product related to the cubic Gauss sum, III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):310–324, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matui:2001:EOC

[70] Hiroki Matui. Ext and orderext classes of certain automorphisms of C^{∗}-algebras arising from Cantor minimal systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):325–354, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nica:2001:DEF

[71] Alexandru Nica, Dimitri Shlyakhtenko, and Roland Speicher. R-diagonal elements and freeness with amalgamation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):355–381, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pivato:2001:BSS

[72] Marcus Pivato. Building a stationary stochastic process from a finitedimensional marginal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):382–413, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rivat:2001:NPF

[73] Joël Rivat and Patrick Sargos. Nombres premiers de la forme floor n^{c}. (French) [Prime numbers of the form floor n^{c}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):414–433, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderPoorten:2001:VDE

[74] Alfred J. van der Poorten and Kenneth S. Williams. Values of the Dedekind eta function at quadratic irrationalities: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):

434–448, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [4].

Akbary:2001:DRP

[75] Amir Akbary and V. Kumar Murty. Descending rational points on elliptic curves to smaller fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):449–469, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bauschke:2001:HPC

[76] Heinz H. Bauschke, Osman Güler, Adrian S. Lewis, and Hristo S. Sendov. Hyperbolic polynomials and convex analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(3):470–488, June 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bojanov:2001:BPL

[77] Borislav D. Bojanov, Werner Haußmann, and Geno P. Nikolov. Bivariate polynomials of least deviation from zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):489–505, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davidson:2001:IDN

[78] Kenneth R. Davidson, David W. Kribs, and Miron E. Shpigel. Isometric dilations of non-commuting finite rank n-tuples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):506–545, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erlijman:2001:MSB

[79] Juliana Erlijman. Multi-sided braid type subfactors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):546–591, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Perera:2001:ISM

[80] Francesc Perera. Ideal structure of multiplier algebras of simple C^{∗}-algebras with real rank zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):592–630, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walters:2001:TNC

[81] Samuel G. Walters. K-theory of noncommutative spheres arising from the Fourier automorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):631–674, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ban:2001:JMP

[82] Dubravka Ban. Jacquet modules of parabolically induced representations and Weyl groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):675–695, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Currie:2001:APA

[83] J. Currie and V. Linek. Avoiding patterns in the Abelian sense. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 696–714, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cushman:2001:DSO

[84] Richard Cushman and Jedrzej Śniatycki. Differential structure of orbit spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):715–755, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [166].

Froese:2001:CUB

[85] Richard Froese. Correction to: “Upper Bounds for the Resonance Counting Function of Schrödinger Operators in Odd Dimensions”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):756–757, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [3].

Goulden:2001:ITF

[86] I. P. Goulden, D. M. Jackson, and F. G. Latour. Inequivalent transitive factorizations into transpositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 758–779, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicolaescu:2001:SWI

[87] Liviu I. Nicolaescu. Seiberg–Witten invariants of lens spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):780–808, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:2001:ATG

[88] Guyan Robertson and Tim Steger. Asymptotic K-theory for groups acting on tA_{2} buildings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):809–833, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Veys:2001:ZFK

[89] Willem Veys. Zeta functions and ‘kontsevich invariants’ on singular varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):834–865, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:2001:IPP

[90] Yifan Yang. Inverse problems for partition functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):866–896, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:2001:SEE

[91] Michael A. Bennett. On some exponential equations of S. S. Pillai. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 897–922, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geramita:2001:DHF

[92] Anthony V. Geramita, Tadahito Harima, and Yong Su Shin. Decompositions of the Hilbert function of a set of points in ¶^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):923–943, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ludwig:2001:RIB

[93] J. Ludwig and C. Molitor-Braun. Représentations irréductibles bornées des groupes de Lie exponentiels. (French) [Bounded irreducible representations of exponential Lie groups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):944–978, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagisa:2001:RAC

[94] Masaru Nagisa, Hiroyuki Osaka, and N. Christopher Phillips. Ranks of algebras of continuous C^{∗}-algebra valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):979–1030, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sampson:2001:CMP

[95] G. Sampson and P. Szeptycki. The complete (L^{p}, L^{p}) mapping properties of some oscillatory integrals in several dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):1031–1056, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:2001:PTL

[96] N. Th. Varopoulos. Potential theory in Lipschitz domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):1057–1120, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Athanasiadis:2001:MPZ

[97] Christos A. Athanasiadis and Francisco Santos. Monotone paths on zonotopes and oriented matroids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 1121–1140, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bushnell:2001:CPT

[98] Colin J. Bushnell and Guy Henniart. Sur le comportement, par torsion, des facteurs epsilon de paires. (French) [Behavior, by twisting, epsilon-pair factors]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53 (??):1141–1173, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Loewen:2001:GVP

[99] Philip D. Loewen and Xianfu Wang. A generalized variational principle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 1174–1193, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Louboutin:2001:EUB

[100] Stéphane Louboutin. Explicit upper bounds for residues of Dedekind zeta functions and values of L-functions at s = 1, and explicit lower bounds for relative class numbers of CM-fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??): 1194–1222, ???? 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mygind:2001:CCS

[101] Jesper Mygind. Classification of certain simple C^{∗}-algebras with torsion in K_{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):1223–1308, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steer:2001:DHK

[102] Brian Steer and Andrew Wren. The Donaldson–Hitchin–Kobayashi correspondence for parabolic bundles over orbifold surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(??):1309–1339, ???? 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2001:AII

[103] Anonymous. Author index — index des auteurs — for 2001 — pour 2001. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(6):1340– 1343, December 2001. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alekseev:2002:QPM

[104] A. Alekseev, Y. Kosmann-Schwarzbach, and E. Meinrenken. Quasi-Poisson manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):3–29, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Treloar:2002:SGP

[105] Thomas Treloar. The symplectic geometry of polygons in the 3-sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):30–54, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ban:2002:MFQ

[106] Chunsheng Ban, Lee J. McEwan, and András Némethi. On the Milnor fiber of a quasi-ordinary surface singularity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):55–70, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:2002:SPS

[107] Kwok-Kwong Stephen Choi and Jianya Liu. Small prime solutions of quadratic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):71–91, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mezo:2002:CGL

[108] Paul Mezo. Comparisons of general linear groups and their metaplectic coverings I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):92–137, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Razak:2002:CSS

[109] Shaloub Razak. On the classification of simple stably projectionless C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):138–224, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arslan:2002:SWF

[110] Bora Arslan, Alexander P. Goncharov, and Mefharet Kocatepe. Spaces of Whitney functions on Cantor-type sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):225–238, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cartwright:2002:ESP

[111] Donald I. Cartwright and Tim Steger. Elementary symmetric polynomials in numbers of modulus 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):239–262, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chaudouard:2002:IOP

[112] Pierre-Henri Chaudouard. Intégrales orbitales pondérées sur les algèbres de Lie: le cas p-adique. (French) [Weighted orbital integrals on Lie algebras: the p-adic case]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):263–302, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ghahramani:2002:CFC

[113] Fereidoun Ghahramani and Sandy Grabiner. Convergence factors and compactness in weighted convolution algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):303–323, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:2002:PRU

[114] Ian Graham, Hidetaka Hamada, and Gabriela Kohr. Parametric representation of univalent mappings in several complex variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):324–351, ????

2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haines:2002:CCS

[115] Thomas J. Haines. On connected components of Shimura varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 352–395, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lebel:2002:FSS

[116] André Lebel. Framed stratified sets in Morse theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):396–416, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wooley:2002:SES

[117] Trevor D. Wooley. Slim exceptional sets for sums of cubes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):417–448, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akrout:2002:TVE

[118] H. Akrout. Théorème de Voronoí dans les espaces symétriques. (French) [Voronoí theorem in symmetric spaces]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):449–467, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:2002:MMD

[119] David W. Boyd and Fernando Rodriguez-Villegas. Mahler’s measure and the dilogarithm (I). Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):468–492, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Braden:2002:PSG

[120] Tom Braden. Perverse sheaves on Grassmannians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):493–532, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Castelle:2002:AFP

[121] Nathalie Castelle. Approximations fortes pour des processus bivariés. (French) [Strong approximations for bivariate processes]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):533–553, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hausen:2002:EES

[122] Jürgen Hausen. Equivariant embeddings into smooth toric varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 554–570, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2002:DPD

[123] Chi-Kwong Li and Yiu-Tung Poon. Diagonals and partial diagonals of sum of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):571–594, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nahlus:2002:LAP

[124] Nazih Nahlus. Lie algebras of proaffine algebraic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):595–607, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stanley:2002:LSC

[125] Donald Stanley. On the Lusternik– Schnirelmann category of maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 608–633, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weber:2002:FSW

[126] Eric Weber. Frames and single wavelets for unitary groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):634–647, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yuan:2002:RSP

[127] Wenjun Yuan and Yezhou Li. Rational solutions of Painlevé equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 648–672, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asgari:2002:LFS

[128] Mahdi Asgari. Local L-functions for split spinor groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):673–693, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gabriel:2002:CAS

[129] Michael J. Gabriel. Cuntz algebra states defined by implementers of endomorphisms of the CAR algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 694–708, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:2002:IMR

[130] Mourad E. H. Ismail and Dennis Stanton. q-integral and moment representations for q-orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):709–735, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kearnes:2002:CFS

[131] K. A. Kearnes, E. W. Kiss, Á. Szendrei, and R. D. Willard. Chief factor sizes in finitely generated varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 736–756, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larose:2002:SPG

[132] Benoit Larose. Strongly projective graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):757–768, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miyazaki:2002:NOW

[133] Takuya Miyazaki. Nilpotent orbits and Whittaker functions for derived functor modules of Sp(2, R). Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):769–794, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moller:2002:STT

[134] Rögnvaldur G. Möller. Structure theory of totally disconnected locally compact groups via graphs and permutations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):795–827, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moriyama:2002:SFS

[135] Tomonori Moriyama. Spherical functions for the semisimple symmetric pair (Sp(2, R), SL(2, C)). Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):828–896, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayuso:2002:VTF

[136] Pedro Fortuny Ayuso. The valuative theory of foliations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):897–915, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bastien:2002:CCM

[137] G. Bastien and M. Rogalski. Convexité, complète monotonie et inégalités sur les fonctions zêta et gamma sur les fonctions des opérateurs de Baskakov et sur des fonctions arithmétiques. (French) [convexity, complete monotonicity, and inequality for zeta functions and gamma functions of the Baskakov operators and for arithmetic functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(5): 916–944, October 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boivin:2002:ACS

[138] André Boivin, Paul M. Gauthier, and Petr V. Paramonov. Approximation on closed sets by analytic or meromorphic solutions of elliptic equations and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):945–969, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cegarra:2002:GCG

[139] A. M. Cegarra, J. M. García-Calcines, and J. A. Ortega. On graded categorical groups and equivariant group extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):970–997, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dimassi:2002:RSV

[140] Mouez Dimassi. Resonances for slowly varying perturbations of a periodic Schrödinger operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):998–1037, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gavrilov:2002:BLC

[141] Lubomir Gavrilov and Iliya D. Iliev. Bifurcations of limit cycles from infinity in quadratic systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1038–1064, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayashi:2002:LTB

[142] Nakao Hayashi and Pavel I. Naumkin. Large time behavior for the cubic nonlinear Schrödinger equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1065–1085, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Polterovich:2002:CHT

[143] Iosif Polterovich. Combinatorics of the heat trace on spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1086–1099, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wood:2002:OBF

[144] Peter J. Wood. The operator biprojectivity of the Fourier algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1100–1120, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bao:2002:FNE

[145] Jiguang Bao. Fully nonlinear elliptic equations on general domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1121–1141, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binding:2002:FDE

[146] Paul Binding and Branko Ćurgus. Form domains and eigenfunction expansions for differential equations with eigenparameter dependent boundary conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1142–1164, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blasco:2002:MVV

[147] Oscar Blasco and José Luis Arregui. Multipliers on vector valued Bergman spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1165–1186, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cobo:2002:IMR

[148] Milton Cobo, Carlos Gutierrez, and Jaume Llibre. On the injectivity of C^{1} maps of the real plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1187–1201, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fernandez:2002:OGR

[149] J. Fernández, J-C. Lario, and A. Rio. Octahedral Galois representations arising from Q-curves of degree 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1202–1228, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gow:2002:WCU

[150] Roderick Gow and Fernando Szechtman. The Weil character of the unitary group associated to a finite local ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1229–1253, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaev:2002:EAU

[151] A. V. Isaev and N. G. Kruzhilin. Effective actions of the unitary group on complex manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1254–1279, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Skrzypczak:2002:BSH

[152] Leszek Skrzypczak. Besov spaces and Hausdorff dimension for some Carnot– Carathéodory metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1280–1304, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vulakh:2002:CFA

[153] L. Ya. Vulakh. Continued fractions associated with SL_{3}(Z) and units in complex cubic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??):1305–1318, ???? 2002. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yekutieli:2002:CHC

[154] Amnon Yekutieli. The continuous Hochschild cochain complex of a scheme. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54 (??):1319–1337, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2002:AII

[155] Anonymous. Author index — index des auteurs — for 2002 — pour 2002. Canadian Journal of Mathematics = Journal canadien de mathématiques, 54(??): 1338–1342, ???? 2002. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baake:2003:ESM

[156] Michael Baake and Ellen Baake. An exactly solved model for mutation, recombination and selection. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):3–41, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [434].

Benanti:2003:SVG

[157] Francesca Benanti, Onofrio M. Di Vincenzo, and Vincenzo Nardozza. ∗- subvarieties of the variety generated by (M_{2}(K), t). Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):42–63, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Braun:2003:HOT

[158] Rüdiger W. Braun, Reinhold Meise, and B. A. Taylor. Higher order tangents to analytic varieties along curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):64–90, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:2003:SCF

[159] Man-Duen Choi, Chi-Kwong Li, and Yiu-Tung Poon. Some convexity features associated with unitary orbits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):91–111, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shen:2003:FM

[160] Zhongmin Shen. Finsler metrics with K = 0 and S = 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):112–132, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shimada:2003:ZVK

[161] Ichiro Shimada. On the zariski-van Kampen theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):133–156, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shimada:2003:ZHS

[162] Ichiro Shimada. Zariski hyperplane section theorem for Grassmannian varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):157–180, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Theriault:2003:HDI

[163] Stephen D. Theriault. Homotopy decompositions involving the loops of coassociative co-H spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):181–203, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yan:2003:NCO

[164] Yaqiang Yan. On the nonsquare constants of Orlicz spaces with Orlicz norm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):204–224, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banks:2003:SKS

[165] William D. Banks, Asma Harcharras, and Igor E. Shparlinski. Short Kloosterman sums for polynomials over finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):225–246, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cushman:2003:DSO

[166] Richard Cushman and Jedrzej Śniatycki. “Differential Structure of Orbit Spaces”: Erratum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(2):247, April 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [84].

Dhillon:2003:GTT

[167] Ajneet Dhillon. A generalized Torelli theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):248–265, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kogan:2003:TAM

[168] Irina A. Kogan. Two algorithms for a moving frame construction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 266–291, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pitman:2003:IDL

[169] Jim Pitman and Marc Yor. Infinitely divisible laws associated with hyperbolic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):292–330, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Savitt:2003:MNP

[170] David Savitt. The maximum number of points on a curve of genus 4 over F_{8} is 25. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):331–352, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Silberger:2003:WEM

[171] Allan J. Silberger and Ernst-Wilhelm Zink. Weak explicit matching for level zero discrete series of unit groups of √-

adic simple algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):353–378, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stessin:2003:GFH

[172] Michael Stessin and Kehe Zhu. Generalized factorization in Hardy spaces and the commutant of Toeplitz operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):379–400, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varopoulos:2003:GEL

[173] N. Th. Varopoulos. Gaussian estimates in Lipschitz domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):401–431, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zaharescu:2003:PCS

[174] Alexandru Zaharescu. Pair correlation of squares in p-adic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):

432–448, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Albeverio:2003:GSS

[175] Sergio Albeverio, Konstantin A. Makarov, and Alexander K. Motovilov. Graph subspaces and the spectral shift function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):449–503, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2003:COR

[176] Jiecheng Chen, Dashan Fan, and Yiming Ying. Certain operators with rough singular kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):504–532, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edo:2003:AME

[177] Eric Edo. Automorphismes modérés de l’espace affine. (French) [Moderate automorphisms of affine space]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 533–560, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laface:2003:QHL

[178] Antonio Laface and Luca Ugaglia. Quasi-homogeneous linear systems on P^{2} with base points of multiplicity 5. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):561–575, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lukashov:2003:AOE

[179] A. L. Lukashov and F. Peherstorfer. Automorphic orthogonal and extremal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):576–608, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moraru:2003:ISA

[180] Ruxandra Moraru. Integrable systems associated to a Hopf surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 609–635, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwartzman:2003:HDA

[181] Sol Schwartzman. Higher dimensional asymptotic cycles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):636–648, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zucconi:2003:SIP

[182] Francesco Zucconi. Surfaces with p_{g} = q = 2 and an irrational pencil. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 649–672, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:2003:NCE

[183] Greg W. Anderson and Yi Ouyang. A note on cyclotomic Euler systems and the double complex method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):

673–692, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borne:2003:FRR

[184] Niels Borne. Une formule de Riemann– Roch équivariante pour les courbes. (French) [A formula of Riemann–Roch for equivariant curves]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):693–710, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Broughan:2003:ATR

[185] Kevin A. Broughan. Adic topologies for the rational integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):711–723, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cao:2003:SLP

[186] Xifang Cao, Qingkai Kong, Hongyou Wu, and Anton Zettl. Sturm–Liouville problems whose leading coefficient function changes sign. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):724–749, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gobel:2003:AFR

[187] Rüdiger Göbel, Saharon Shelah, and Lutz Strüngmann. Almost-free E-rings of cardinality ℵ_{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):750–765, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kerler:2003:HTA

[188] Thomas Kerler. Homology TQFT’s and the Alexander–Reidemeister invariant of 3-manifolds via Hopf algebras and skein theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):766–821, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:2003:OGP

[189] Djun Maximilian Kim and Dale Rolfsen.

An ordering for groups of pure braids and fibre-type hyperplane arrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):822–838, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2003:CCT

[190] Min Ho Lee. Cohomology of complex torus bundles associated to cocycles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):839–855, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Su:2003:PBS

[191] Yucai Su. Poisson brackets and structure of nongraded Hamiltonian Lie algebras related to locally-finite derivations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):856–896, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Archinard:2003:HAV

[192] Natália Archinard. Hypergeometric Abelian varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):897–932, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:2003:RP

[193] Jennifer Beineke and Daniel Bump. Renormalized periods on GL(3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 933–968, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glockner:2003:LGM

[194] Helge Glöckner. Lie groups of measurable mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):969–999, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graczyk:2003:SCR

[195] P. Graczyk and P. Sawyer. Some convexity results for the Cartan decomposition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1000–1018, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handelman:2003:MEP

[196] David Handelman. More eventual positivity for analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 1019–1079, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kellerhals:2003:QSG

[197] Ruth Kellerhals. Quaternions and some global properties of hyperbolic 5-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1080–1099, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khesin:2003:PH

[198] Boris Khesin and Alexei Rosly. Polar homology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1100–1120, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bettaieb:2003:CRT

[199] Karem Bettaïeb. Classification des représentations tempérées d’un groupe p-adique. (French) [Classification of representations of a temperate padic group]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1121–1133, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casarino:2003:NCH

[200] Valentina Casarino. Norms of complex harmonic projection operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??): 1134–1154, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokovic:2003:CON

[201] Dragomir Ž. – Doković and Michael Litvinov. The closure ordering of nilpotent orbits of the complex symmetric pair (SO_{p+q}, SO_{p} × SO_{q}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1155–1190, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Granville:2003:DMV

[202] Andrew Granville and K. Soundararajan. Decay of mean values of multiplicative functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1191–1230, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Havin:2003:AMMa

[203] Victor Havin and Javad Mashreghi. Admissible majorants for model subspaces of H^{2}, Part I: Slow winding of the generating inner function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1231–1263, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Havin:2003:AMMb

[204] Victor Havin and Javad Mashreghi. Admissible majorants for model subspaces of H^{2}, Part II: Fast winding of the generating inner function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):1264–1301, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katsura:2003:ISC

[205] Takeshi Katsura. The ideal structures of crossed products of Cuntz algebras by quasi-free actions of Abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55 (??):1302–1338, ???? 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2003:AII

[206] Anonymous. Author index — index des auteurs — for 2003 — pour 2003. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(6):1339– 1342, December 2003. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amini:2004:LCP

[207] Massoud Amini. Locally compact pro-C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):3–22, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:2004:TDE

[208] Michael A. Bennett and Chris M.

Skinner. Ternary Diophantine equations via Galois representations and modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):23–54, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harper:2004:E

[209] Malcolm Harper. Z[\sqrt{14}] is Euclidean.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):55–70, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harper:2004:ERA

[210] Malcolm Harper and M. Ram Murty.

Euclidean rings of algebraic integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):71–76, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holmes:2004:HDG

[211] Mark Holmes, Antal A. Járai, Akira Sakai, and Gordon Slade. Highdimensional graphical networks of selfavoiding walks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):77–114, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kenny:2004:EHD

[212] Robert Kenny. Estimates of Hausdorff dimension for the non-wandering set of an open planar billiard. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):115–133, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2004:LOM

[213] Chi-Kwong Li and Ahmed Ramzi Sourour. Linear operators on matrix algebras that preserve the numerical range, numerical radius or the states. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):134–167, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pogge:2004:CRS

[214] James Todd Pogge. On a certain residual spectrum of Sp_{8}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):168–193, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saikia:2004:SGE

[215] A. Saikia. Selmer groups of elliptic curves with complex multiplication. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):194–208, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmuland:2004:CLT

[216] Byron Schmuland and Wei Sun. A central limit theorem and law of the iterated logarithm for a random field with exponential decay of correlations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 209–224, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blower:2004:CUC

[217] Gordon Blower and Thomas Ransford. Complex uniform convexity and Riesz measure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):225–245, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bonnafe:2004:EUR

[218] Cédric Bonnafé. Éléments unipotents réguliers des sous-groupes de Levi. (French) [Unipotent regular elements of Levi subgroups ]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):246–276, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dostanic:2004:SPC

[219] Milutin R. Dostanić. Spectral properties of the commutator of Bergman’s projection and the operator of multiplication by an analytic function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):277–292, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khomenko:2004:SMI

[220] Oleksandr Khomenko and Volodymyr Mazorchuk. Structure of modules induced from simple modules with minimal annihilator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):293–309, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Llibre:2004:GQD

[221] Jaume Llibre and Dana Schlomiuk. The geometry of quadratic differential systems with a weak focus of third order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):310–343, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miao:2004:PMA

[222] Tianxuan Miao. Predual of the multiplier algebra of A_{p}(G) and amenability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):344–355, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murty:2004:NAG

[223] M. Ram Murty and Filip Saidak. Non-Abelian generalizations of the Erdős– Kac theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):356–372, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Orton:2004:EPW

[224] Louisa Orton. An elementary proof of a weak exceptional zero conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):

373–405, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pal:2004:TSE

[225] Ambrus Pál. Theta series, Eisenstein series and Poincaré series over function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):406–430, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenblatt:2004:GAS

[226] Joseph Rosenblatt and Michael Taylor. Group actions and singular martingales II, the recognition problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 431–448, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Demeter:2004:BCA

[227] Ciprian Demeter. The best constants associated with some weak maximal inequalities in ergodic theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 449–471, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fonf:2004:IDP

[228] Vladimir P. Fonf and Libor Veselý.

Infinite-dimensional polyhedrality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 472–494, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gomi:2004:CAF

[229] Yasushi Gomi, Iku Nakamura, and Ken ichi Shinoda. Coinvariant algebras of finite subgroups of SL(3, C). Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 495–528, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martinez-Finkelshtein:2004:AMD

[230] A. Martínez-Finkelshtein, V. Maymeskul, E. A. Rakhmanov, and E. B. Saff. Asymptotics for minimal discrete Riesz energy on curves in R^{d}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):529–552, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohammadalikhani:2004:CRS

[231] Ramin Mohammadalikhani. Cohomology ring of symplectic quotients by circle actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):553–565, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ni:2004:GMH

[232] Yilong Ni. Geodesics in a manifold with Heisenberg group as boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 566–589, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ni:2004:HKG

[233] Yilong Ni. The heat kernel and Green’s function on a manifold with Heisenberg group as boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):590–611, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pal:2004:SPP

[234] Ambrus Pál. Solvable points on projective algebraic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):612–637, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sniatycki:2004:MRP

[235] Jedrzej Śniatycki. Multisymplectic reduction for proper actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 638–654, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tao:2004:NPS

[236] Xiangxing Tao and Henggeng Wang.

On the Neumann problem for the Schrödinger equations with singular potentials in Lipschitz domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 655–672, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cali:2004:DSS

[237] Élie Cali. Défaut de semi-stabilité des courbes elliptiques dans le cas non ramifié. (French) [Semi-stability failure of elliptic curves in the unbranched case]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):673–698, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaspari:2004:BFH

[238] Thierry Gaspari. Bump functions with Hölder derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):699–715, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guardo:2004:FPT

[239] Elena Guardo and Adam Van Tuyl. Fat points in P^{1}×P^{1} and their Hilbert functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):716–741, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jiang:2004:SCC

[240] Chunlan Jiang. Similarity classification of Cowen–Douglas operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 742–775, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lim:2004:BAR

[241] Yongdo Lim. Best approximation in Riemannian geodesic submanifolds of positive definite matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):776–793, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Michel:2004:SCB

[242] Laurent Michel. Semi-classical behavior of the scattering amplitude for trapping perturbations at fixed energy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 794–824, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Penot:2004:DPO

[243] Jean-Paul Penot. Differentiability properties of optimal value functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 825–842, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruan:2004:TDR

[244] Zhong-Jin Ruan. Type decomposition and the rectangular AFD property for W^{∗}-TRO’s. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):843–870, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schocker:2004:LEK

[245] Manfred Schocker. Lie elements and Knuth relations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):871–882, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tandra:2004:KTC

[246] Haryono Tandra and William Moran.

Kirillov theory for a class of discrete nilpotent groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):883–896, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2004:FEA

[247] Jonathan M. Borwein, David Borwein, and William F. Galway. Finding and excluding b-ary Machin-type individual digit formulae. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(5):897–925, October 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hadfield:2004:HRA

[248] Tom Hadfield. K-homology of the rotation algebras A_{θ}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):926–944, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helminck:2004:SQA

[249] Aloysius G. Helminck and Gerald W. Schwarz. Smoothness of quotients associated with a pair of commuting involutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):945–962, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishiwata:2004:BET

[250] Satoshi Ishiwata. A Berry–Esseen type theorem on nilpotent covering graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):963–982, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Junge:2004:FTU

[251] Marius Junge. Fubini’s theorem for ultraproducts of noncommutative L_{p}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):983–1021, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matignon:2004:NOS

[252] D. Matignon and N. Sayari. Nonorientable surfaces and Dehn surgeries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):1022–1033, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rouleux:2004:SCI

[253] Michel Rouleux. Semi-classical integrability,hyperbolic flows and the Birkhoff normal form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1034–1067, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinbach:2004:REG

[254] Anja Steinbach and Hendrik Van Maldeghem. Regular embeddings of generalized hexagons. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1068–1093, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:2004:CLI

[255] Hugh Thomas. Cycle-level intersection theory for toric varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 1094–1120, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chaumat:2004:DPP

[256] Jacques Chaumat and Anne-Marie Chollet. Division par un polynôme hyperbolique. (French) [Division by a hyperbolic polynomial]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1121–1144, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daigle:2004:LHP

[257] Daniel Daigle and Peter Russell. On log Q-homology planes and weighted projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1145–1189, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frank:2004:MFS

[258] Günter Frank, Xinhou Hua, and Rémi Vaillancourt. Meromorphic functions sharing the same zeros and poles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 1190–1227, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ho:2004:CMS

[259] Nan-Kuo Ho and Chiu-Chu Melissa Liu. On the connectedness of moduli spaces of flat connections over compact surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1228–1236, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kishimoto:2004:CSA

[260] Akitaka Kishimoto. Central sequence algebras of a purely infinite simple C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1237–1258, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paterson:2004:FAL

[261] Alan L. T. Paterson. The Fourier algebra for locally compact groupoids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56 (??):1259–1289, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scull:2004:EFA

[262] Laura Scull. Equivariant formality for actions of torus groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??): 1290–1307, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhao:2004:VMH

[263] Jianqiang Zhao. Variations of mixed Hodge structures of multiple polylogarithms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1308–1338, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2004:AII

[264] Anonymous. Author index — index des auteurs — for 2004 — pour 2004. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(6):1339– 1342, December 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alberich-Carraminana:2005:EDA

[265] Maria Alberich-Carramiñana and Joaquim Roé. Enriques diagrams and adjacency of planar curve singularities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):3–16, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bedos:2005:ACA

[266] Erik Bédos, Roberto Conti, and Lars Tuset. On amenability and co-amenability of algebraic quantum groups and their corepresentations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):17–60, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binding:2005:OSS

[267] Paul Binding and Vladimir Strauss. On operators with spectral square but without resolvent points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):61–81, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fallat:2005:JST

[268] Shaun M. Fallat and Michael I. Gekhtman. Jordan structures of totally nonnegative matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):82–98, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fegan:2005:SOO

[269] H. D. Fegan and B. Steer. Second order operators on a compact Lie group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):99–113, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flaschka:2005:BFS

[270] Hermann Flaschka and John Millson. Bending flows for sums of rank one matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):114–158, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jantzen:2005:DSI

[271] Chris Jantzen. Duality and supports of induced representations for orthogonal groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):159–179, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Somodi:2005:SWS

[272] Marius Somodi. On the size of the wild set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):180–203, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xiong:2005:DBC

[273] Jie Xiong and Xiaowen Zhou. On the duality between coalescing Brownian motions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):204–224, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Booss-Bavnbek:2005:UFO

[274] Bernhelm Booss-Bavnbek, Matthias Lesch, and John Phillips. Unbounded Fredholm operators and spectral flow. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):225–250, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cocos:2005:SNR

[275] M. Cocos. Some new results on L^{2} cohomology of negatively curved Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):251–266, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conrad:2005:PEP

[276] Keith Conrad. Partial Euler products on the critical line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):267–297, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kumchev:2005:WGP

[277] Angel V. Kumchev. On the Waring– Goldbach problem: Exceptional sets for sums of cubes and higher powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 298–327, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:2005:CBS

[278] Wentang Kuo and M. Ram Murty. On a conjecture of Birch and Swinnerton-Dyer. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):328–337, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lange:2005:CES

[279] Tanja Lange and Igor E. Shparlinski. Certain exponential sums and random walks on elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):338–350, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:2005:ESA

[280] Huaxin Lin. Extensions by simple C^{∗}-algebras: Quasidiagonal extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):351–399, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sabourin:2005:GC

[281] Sindi Sabourin. Generalized kconfigurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):400–415, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wise:2005:AFP

[282] Daniel T. Wise. Approximating flats by periodic flats in CAT(0) square complexes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):416–448, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alkan:2005:SGF

[283] Emre Alkan. On the sizes of gaps in the Fourier expansion of modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):449–470, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ciesielski:2005:SCS

[284] Krzysztof Ciesielski and Janusz Pawlikowski. Small coverings with smooth functions under the covering property axiom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):471–493, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedlander:2005:SFC

[285] John B. Friedlander and Henryk Iwaniec. Summation formulae for coefficients of L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):494–505, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:2005:RHS

[286] Leonard Gross and Martin Grothaus. Reverse hypercontractivity for subharmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):506–534, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:2005:LFN

[287] Henry H. Kim. On local L-functions and normalized intertwining operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):535–597, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kornelson:2005:LSL

[288] Keri A. Kornelson. Local solvability of Laplacian difference operators arising from the discrete Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):598–615, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muic:2005:RGP

[289] Goran Muić. Reducibility of generalized principal series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):616–647, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nevins:2005:BRP

[290] Monica Nevins. Branching rules for principal series representations of SL(2) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):648–672, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Androulakis:2005:SSM

[291] G. Androulakis, E. Odell, Th. Schlumprecht, and N. Tomczak-Jaegermann. On the structure of the spreading models of a Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):673–707, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finster:2005:CEA

[292] Felix Finster and Margarita Kraus. Curvature estimates in asymptotically flat Lorentzian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):708–723, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Purnaprajna:2005:SRS

[293] B. P. Purnaprajna. Some results on surfaces of general type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):724–749, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sabourin:2005:STO

[294] Hervé Sabourin. Sur la structure transverse à une orbite nilpotente adjointe. (French) [On the transverse structure of a nilpotent adjoint orbit]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 750–770, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schrohe:2005:RCE

[295] E. Schrohe and J. Seiler. The resolvent of closed extensions of cone differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):771–811, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Trifkovic:2005:VIE

[296] Mak Trifković. On the vanishing of μ-invariants of elliptic curves over Q. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (4):812–843, August 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:2005:PS

[297] Gordon Williams. Petrie schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):

844–870, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2005:HYM

[298] Xi Zhang. Hermitian Yang–Mills–Higgs metrics on complete Kähler manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):871–896, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berezhnoi:2005:RBI

[299] Evgenii I. Berezhnoi and Lech Maligranda. Representation of Banach ideal spaces and factorization of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):897–940, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:2005:STH

[300] Christian Berg and Antonio J. Durán.

Some transformations of Hausdorff moment sequences and harmonic numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):941–960, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2005:CMF

[301] Jonathan M. Borwein and Xianfu Wang.

Cone-monotone functions: Differentiability and continuity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):961–982, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

anHuef:2005:SIT

[302] Astrid an Huef, Iain Raeburn, and Dana P. Williams. A symmetric imprimitivity theorem for commuting proper actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):983–1011, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karigiannis:2005:DS

[303] Spiro Karigiannis. Deformations of G_{2} and Spin(7) structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 1012–1055, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ozawa:2005:HGA

[304] Narutaka Ozawa and Marc A. Rieffel. Hyperbolic group C^{∗}-algebras and freeproduct C^{∗}-algebras as compact quantum metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1056–1079, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pritsker:2005:GSM

[305] Igor E. Pritsker. The Gelfond– Schnirelman method in prime number theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1080–1101, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weston:2005:PRF

[306] Tom Weston. Power residues of Fourier coefficients of modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):

1102–1120, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barr:2005:EEA

[307] Michael Barr, R. Raphael, and R. G.

Woods. On CR-epic embeddings and absolute CR-epic spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1121– 1138, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burke:2005:MWE

[308] Maxim R. Burke and Arnold W. Miller.

Models in which every nonmeager set is nonmeager in a nowhere dense Cantor set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):1139–1154, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cojocaru:2005:SSL

[309] Alina Carmen Cojocaru, Etienne Fouvry, and M. Ram Murty. The square sieve and the Lang–Trotter conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57 (??):1155–1177, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cutkosky:2005:ABL

[310] Steven Dale Cutkosky, Huy Tài Hà, Hema Srinivasan, and Emanoil Theodorescu. Asymptotic behavior of the length of local cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 1178–1192, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dungey:2005:SCD

[311] Nick Dungey. Some conditions for decay of convolution powers and heat kernels on groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1193–1214, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khare:2005:RLC

[312] Chandrashekhar Khare. Reciprocity law for compatible systems of Abelian modp Galois representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 1215–1223, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kopotun:2005:CPA

[313] K. A. Kopotun, D. Leviatan, and I. A. Shevchuk. Convex polynomial approximation in the uniform norm: Conclusion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1224–1248, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindstrom:2005:SSC

[314] Mikael Lindström, Eero Saksman, and Hans-Olav Tylli. Strictly singular and cosingular multiplications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1249– 1278, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maad:2005:SPH

[315] Sara Maad. A semilinear problem for the Heisenberg Laplacian on unbounded domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1279–1290, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Riveros:2005:DH

[316] Carlos M. C. Riveros and Keti Tenenblat. Dupin hypersurfaces in R^{5}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 1291–1313, ???? 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhitomirskii:2005:RDT

[317] M. Zhitomirskii. Relative Darboux theorem for singular manifolds and local contact algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1314–1340, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2005:AII

[318] Anonymous. Author index — index des auteurs — for 2005 — pour 2005. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(6):1341– 1344, December 2005. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Said:2006:FEZ

[319] Salem Ben Saïd. The functional equation of zeta distributions associated with non-Euclidean Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):

3–22, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dabbaghian-Abdoly:2006:CRF

[320] Vahid Dabbaghian-Abdoly. Constructing representations of finite simple groups and covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):23–38, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Exel:2006:AID

[321] R. Exel and A. Vershik. C^{∗}-algebras of irreversible dynamical systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 39–63, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Filippakis:2006:MRN

[322] Michael Filippakis, Leszek Gasiński, and Nikolaos S. Papageorgiou. Multiplicity results for nonlinear Neumann problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):64–92, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:2006:MHM

[323] Julia Gordon. Motivic Haar measure on reductive groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):93–114, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ivorra:2006:QRE

[324] W. Ivorra and A. Kraus. Quelques résultats sur les équations ax^{p} + by^{p} = cz^{2}. (French) [Some results for the equations ax^{p} + by^{p} = cz^{2}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):115–153, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prestini:2006:SIP

[325] Elena Prestini. Singular integrals on product spaces related to the Carleson operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):154–179, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reiten:2006:IDR

[326] Idun Reiten and Claus Michael Ringel.

Infinite dimensional representations of canonical algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):180–224, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Azam:2006:GRL

[327] Saeid Azam. Generalized reductive Lie algebras: Connections with extended affine Lie algebras and Lie tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 225–248, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hernandez:2006:CFP

[328] M. Bello Hernández and J. Mínguez Ceniceros. Convergence of Fourier– Padé approximants for Stieltjes functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):249–261, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biswas:2006:CPP

[329] Indranil Biswas. Connections on a parabolic principal bundle over a curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (??):262–281, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fels:2006:NRH

[330] M. E. Fels and A. G. Renner. Non-reductive homogeneous pseudo-Riemannian manifolds of dimension four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):282–311, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gamblin:2006:PIR

[331] Didier Gamblin. Partie imaginaire des résonances de Rayleigh dans le cas d’une boule. (French) [Imaginary part of Rayleigh resonances in the case of a ball]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):312–343, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldberg:2006:RGE

[332] David Goldberg. Reducibility for SU_{n} and generic elliptic representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (??):344–361, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldin:2006:CPS

[333] R. F. Goldin and S. Martin. Cohomology pairings on the symplectic reduction of products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):362–380, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jakobson:2006:EMF

[334] Dmitry Jakobson, Nikolai Nadirashvili, and Iosif Polterovich. Extremal metric for the first eigenvalue on a Klein bottle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (??):381–400, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kolountzakis:2006:PEP

[335] Mihail N. Kolountzakis and Szilárd Gy.

Révész. On pointwise estimates of positive definite functions with given support. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):401–418, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snaith:2006:SCN

[336] Victor P. Snaith. Stark’s conjecture and new Stickelberger phenomena. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 419–448, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agarwal:2006:EMP

[337] Ravi P. Agarwal, Daomin Cao, Haishen Lü, and Donal O’Regan. Existence and multiplicity of positive solutions for singular semipositone p-Laplacian equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):449–475, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chipalkatti:2006:ASA

[338] Jaydeep Chipalkatti. Apolar schemes of algebraic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):476–491, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chua:2006:ETW

[339] Seng-Kee Chua. Extension theorems on weighted Sobolev spaces and some applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):492–528, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dijkstra:2006:GHR

[340] Jan J. Dijkstra and Jan van Mill. On the group of homeomorphisms of the real line that map the pseudoboundary onto itself. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):529–547, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Galanopoulos:2006:HQH

[341] P. Galanopoulos and M. Papadimitrakis. Hausdorff and quasi-Hausdorff matrices on spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 548–579, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greither:2006:ACG

[342] Cornelius Greither and Radan Kucera.

Annihilators for the class group of a cyclic field of prime power degree, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (??):580–599, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martinez-Maure:2006:GSM

[343] Yves Martinez-Maure. Geometric study of Minkowski differences of plane convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):600–624, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohrdieck:2006:SCS

[344] Stephan Mohrdieck. A Steinberg cross section for non-connected affine Kac– Moody groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):625–642, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yu:2006:CTC

[345] Xiaoxiang Yu. Centralizers and twisted centralizers: Application to intertwining operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):643–672, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bart:2006:GCC

[346] Anneke Bart and Kevin P. Scannell. The generalized cuspidal cohomology problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):673–690, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bendikov:2006:HBI

[347] A. Bendikov and L. Saloff-Coste. Hypoelliptic bi-invariant Laplacians on infinite dimensional compact groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 691–725, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chiang:2006:VDT

[348] Yik-Man Chiang and Mourad E. H. Ismail. On value distribution theory of second order periodic ODEs, special functions and orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (4):726–767, August 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [549].

Hu:2006:DNA

[349] Zhiguo Hu and Matthias Neufang. Decomposability of von Neumann algebras and the Mazur property of higher level. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):768–795, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Im:2006:MWG

[350] Bo-Hae Im. Mordell–Weil groups and the rank of elliptic curves over large fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):796–819, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moreno:2006:DMC

[351] J. P. Moreno, P. L. Papini, and R. R.

Phelps. Diametrically maximal and constant width sets in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 820–842, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ozluk:2006:OLD

[352] A. E. Õzlük and C. Snyder. On the onelevel density conjecture for quadratic Dirichlet L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):843–858, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:2006:NIN

[353] C. J. Read. Nonstandard ideals from nonstandard dual pairs for L^{1}(ω) and l^{1}(ω). Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):859–876, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Selick:2006:FDL

[354] P. Selick, S. Theriault, and J. Wu. Functorial decompositions of looped coassociative co-H spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):877–896, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Courtes:2006:DIG

[355] François Courtès. Distributions invariantes sur les groupes réductifs quasi-déployés. (French) [Invariant distributions on quasi-deployed reductive groups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):897–999, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dhillon:2006:CMV

[356] Ajneet Dhillon. On the cohomology of moduli of vector bundles and the Tamagawa number of SL_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1000–1025, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handelman:2006:KRL

[357] David Handelman. Karamata renewed and local limit results. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1026– 1094, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sakellaridis:2006:CSF

[358] Yiannis Sakellaridis. A Casselman– Shalika formula for the Shalika model of GL_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1095–1120, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bownik:2006:FCW

[359] Marcin Bownik and Darrin Speegle. The Feichtinger conjecture for wavelet frames, Gabor frames and frames of translates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1121–1143, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hamana:2006:PAN

[360] Masamichi Hamana. Partial ∗- automorphisms, normalizers, and submodules in monotone complete C^{∗}- algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1144–1202, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heiermann:2006:OUP

[361] Volker Heiermann. Orbites unipotentes et pôles d’ordre maximal de la fonction μ de Harish-Chandra. (French) [Unipotent orbits and poles of maximal order of the Harish-Chandra μ function]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 1203–1228, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henniart:2006:IOT

[362] Guy Henniart and Bertrand Lemaire.

Intégrales orbitales tordues sur GL(n, F) et corps locaux proches: applications. (French) [Twisted orbital integrals on GL(n, F) and close local bodies: applications]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1229–1267, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sims:2006:GII

[363] Aidan Sims. Gauge-invariant ideals in the C^{∗}-algebras of finitely aligned higher-rank graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??):1268–1290, ???? 2006. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weimar-Woods:2006:GSG

[364] Evelyn Weimar-Woods. The general structure of G-graded contractions of Lie algebras I. the classification. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(??): 1291–1340, ???? 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2006:AII

[365] Anonymous. Author index — index des auteurs — for 2006 — pour 2006. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58(6):1341– 1344, December 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biller:2007:HGC

[366] Harald Biller. Holomorphic generation of continuous inverse algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 3–35, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Develin:2007:CDS

[367] Mike Develin, Jeremy L. Martin, and Victor Reiner. Classification of Ding’s Schubert varieties: Finer rook equivalence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):36–62, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferenczi:2007:SRS

[368] Valentin Ferenczi and Elói Medina Galego. Some results on the Schroeder– Bernstein property for separable Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):63–84, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foster:2007:CCN

[369] J. H. Foster and Monika Serbinowska.

On the convergence of a class of nearly alternating series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):85–108, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jayanthan:2007:FCP

[370] A. V. Jayanthan, Tony J. Puthenpurakal, and J. K. Verma. On fiber cones of ⇕-primary ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):109–126, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamzouri:2007:SVI

[371] Youness Lamzouri. Smooth values of the iterates of the Euler phi-function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 127–147, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muic:2007:CCU

[372] Goran Muić. On certain classes of unitary representations for split classical groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):148–185, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Okoh:2007:EAK

[373] F. Okoh and F. Zorzitto. Endomorphism algebras of Kronecker modules regulated by quadratic function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 186–210, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roy:2007:TEA

[374] Damien Roy. On two exponents of approximation related to a real number and its square. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):211–224, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:2007:HAM

[375] Matt Baker and Robert Rumely. Harmonic analysis on metrized graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):225–275, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernardis:2007:WIH

[376] A. L. Bernardis, F. J. Martín-Reyes, and P. Ortega Salvador. Weighted inequalities for Hardy–Steklov operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 276–295, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chein:2007:BLN

[377] Orin Chein and Edgar G. Goodaire. Bol loops of nilpotence class two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):

296–310, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Christianson:2007:GZZ

[378] Hans Christianson. Growth and zeros of the zeta function for hyperbolic rational maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):311–331, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leuschke:2007:ERF

[379] Graham J. Leuschke. Endomorphism rings of finite global dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 332–342, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:2007:WSP

[380] Huaxin Lin. Weak semiprojectivity in purely infinite simple C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):343–371, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maisner:2007:ZFS

[381] Daniel Maisner and Enric Nart. Zeta functions of supersingular curves of genus 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):372–392, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Servat:2007:SPO

[382] E. Servat. Le splitting pour l’opérateur de Klein–Gordon: une approche heuristique et numérique Harish-Chandra. (French) [Splitting for the Klein–Gordon operator: a heuristic numerical Harish-Chandra approach]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):393–417, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stoimenow:2007:CKV

[383] A. Stoimenow. On cabled knots and Vassiliev invariants (not) contained in knot polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):418–448, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Badulescu:2007:ORT

[384] Alexandru Ioan Badulescu. SL_{n}, orthogonality relations and transfer. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 449–464, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barr:2007:SAE

[385] Michael Barr, John F. Kennison, and R. Raphael. Searching for absolute CR-epic spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):465–487, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernardi:2007:OVV

[386] A. Bernardi, M. V. Catalisano, A. Gimigliano, and M. Idà. Osculating varieties of Veronese varieties and their higher secant varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):488–502, ????

2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chevallier:2007:CGT

[387] Nicolas Chevallier. Cyclic groups and the three distance theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 503–552, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dasgupta:2007:CEU

[388] Samit Dasgupta. Computations of elliptic units for real quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 553–574, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hernandez-Hernandez:2007:CIA

[389] Fernando Hernández-Hernández and Michael Hrusák. Cardinal invariants of analytic P -ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):575–595, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Itza-Ortiz:2007:ETM

[390] Benjamín A. Itzá-Ortiz. Eigenvalues, K-theory and minimal flows. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 596–613, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labuschagne:2007:PNO

[391] C. C. A. Labuschagne. Preduals and nuclear operators associated with bounded, p-convex, p-concave and positive psumming operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):614–637, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacDonald:2007:DIN

[392] Gordon W. MacDonald. Distance from idempotents to nilpotents. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 638–657, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minac:2007:DAP

[393] J. Minác and A. Wadsworth. Division algebras of prime degree and maximal Galois p-extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):658–672, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ash:2007:HFD

[394] Avner Ash and Solomon Friedberg. Hecke L-functions and the distribution of totally positive integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 673–695, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bangoura:2007:ALH

[395] Momo Bangoura. Algèbres de Lie d’homotopie associées à une protobigèbre de Lie. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):696–711, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Billig:2007:JM

[396] Yuly Billig. Jet modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 712–729, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdelyi:2007:LSI

[397] T. Erdélyi and D. S. Lubinsky. Large sieve inequalities via subharmonic methods and the Mahler measure of the Fekete polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):730–741, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gil:2007:GSC

[398] Juan B. Gil, Thomas Krainer, and Gerardo A. Mendoza. Geometry and spectra of closed extensions of elliptic cone operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):742–794, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jaworski:2007:CDE

[399] Wojciech Jaworski and Matthias Neufang. The Choquet–Deny equation in a Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):795–827, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ortner:2007:NBR

[400] Ronald Ortner and Wolfgang Woess.

Non-backtracking random walks and cogrowth of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):828–844, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schaffhauser:2007:RFG

[401] Florent Schaffhauser. Representations of the fundamental group of an Lpunctured sphere generated by products of Lagrangian involutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 845–879, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

denvan:2007:RIV

[402] John E. den van. Radical ideals in valuation domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):880–896, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruneau:2007:GSP

[403] Laurent Bruneau. The ground state problem for a quantum Hamiltonian model describing friction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):897–916, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Currey:2007:ACQ

[404] Bradley N. Currey. Admissibility for a class of quasiregular representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):917–942, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finster:2007:WEW

[405] Felix Finster and Margarita Kraus. A weighted L^{2}-estimate of the Witten spinor in asymptotically Schwarzschild manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):943–965, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forrest:2007:OAF

[406] Brian E. Forrest, Volker Runde, and Nico Spronk. Operator amenability of the Fourier algebra in the cbmultiplier norm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(5):966–980, October 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jiang:2007:CRC

[407] Yunfeng Jiang. The Chen–Ruan cohomology of weighted projective spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):981–1007, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaczynski:2007:IZT

[408] Tomasz Kaczynski, Marian Mrozek, and Anik Trahan. Ideas from Zariski topology in the study of cubical homology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):1008–1028, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalton:2007:G

[409] N. J. Kalton, A. Koldobsky, V. Yaskin, and M. Yaskina. The geometry of L_{0}.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):1029–1068, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reydy:2007:QJA

[410] Carine Reydy. Quotients jacobiens: une approche algébrique. (French) [Jacobian quotients: an algebraic approach]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):1069–1097, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rodrigues:2007:RES

[411] B. Rodrigues. Ruled exceptional surfaces and the poles of Motivic zeta functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1098–1120, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alayont:2007:MCS

[412] Feryâl Alayont. Meromorphic continuation of spherical cuspidal data Eisenstein series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1121–1134, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bjorn:2007:SEH

[413] Anders Björn, Jana Björn, and Nageswari Shanmugalingam. Sobolev extensions of Hölder continuous and characteristic functions on metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1135–1153, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boardman:2007:TFS

[414] J. Michael Boardman and W. Stephen Wilson. k(n)-torsion-free H-spaces and P (n)-cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1154–1206, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bu:2007:MRO

[415] Shangquan Bu and Christian Merdy Le. H^{p}-maximal regularity and operator valued multipliers on Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):1207–1222, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buraczewski:2007:CZO

[416] Dariusz Buraczewski, Teresa Martinez, and José L. Torrea. Calderón–Zygmund operators associated to ultraspherical expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1223–1244, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2007:GPI

[417] Qun Chen and Zhen-Rong Zhou. On gap properties and instabilities of p-Yang–Mills fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1245–1259, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deng:2007:GEC

[418] Bangming Deng, Jie Du, and Jie Xiao. Generic extensions and canonical bases for cyclic quivers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1260–1283, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fukshansky:2007:EWD

[419] Lenny Fukshansky. On effective Witt decomposition and the Cartan– Dieudonné theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1284–1300, ???? 2007. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Furioli:2007:SIW

[420] Giulia Furioli, Camillo Melzi, and Alessandro Veneruso. Strichartz inequalities for the wave equation with the full Laplacian on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59 (??):1301–1322, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginzburg:2007:CJL

[421] David Ginzburg and Erez Lapid. On a conjecture of Jacquet, Lai, and Rallis: Some exceptional cases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??): 1323–1340, ???? 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2007:AII

[422] Anonymous. Author index — index des auteurs — for 2007 — pour 2007. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(6):1341– 1344, December 2007. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boroczky:2008:CBM

[423] Károly Böröczky, Károly J. Böröczky, Carsten Schütt, and Gergely Wintsche. Convex bodies of minimal volume, surface area and mean width with respect to thin shells. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):3–32, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Braun:2008:HOT

[424] Rüdiger W. Braun, Reinhold Meise, and B. A. Taylor. Higher order tangents to analytic varieties along curves. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):33–63, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daigle:2008:CLW

[425] Daniel Daigle. Classification of linear weighted graphs up to blowing-up and blowing-down. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):64–87, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Diwadkar:2008:NCC

[426] Jyotsna Mainkar Diwadkar. Nilpotent conjugacy classes in p-adic Lie algebras: The odd orthogonal case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 88–108, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gurjar:2008:ALA

[427] R. V. Gurjar, K. Masuda, M. Miyanishi, and P. Russell. Affine lines on affine surfaces and the Makar–Limanov invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):109–139, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kedlaya:2008:GTC

[428] Kiran S. Kedlaya. On the geometry of ptypical covers in characteristic p. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 140–163, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2008:BSH

[429] Sangyop Lee and Masakazu Teragaito. Boundary structure of hyperbolic 3-manifolds admitting annular and toroidal fillings at large distance. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 164–188, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:2008:FTA

[430] Huaxin Lin. Furstenberg transformations and approximate conjugacy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 189–207, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramakrishna:2008:CGR

[431] Ravi Ramakrishna. Constructing Galois representations with very large image. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):208–221, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Silipo:2008:ASE

[432] James Silipo. Amibes de sommes d’exponentielles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):222–240, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexandrova:2008:SCW

[433] Ivana Alexandrova. Semi-classical wavefront set and Fourier integral operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):241–263, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baake:2008:EES

[434] Michael Baake and Ellen Baake. Erratum to: “An Exactly Solved Model for Recombination, Mutation and Selection”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):264–265, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [156].

Bergeron:2008:ICS

[435] Nantel Bergeron, Christophe Reutenauer, Mercedes Rosas, and Mike Zabrocki. Invariants and coinvariants of the symmetric group in noncommuting variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):266–296, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bini:2008:TFH

[436] G. Bini, I. P. Goulden, and D. M. Jackson. Transitive factorizations in the hyperoctahedral group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):297–312, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:2008:API

[437] Yong-Kab Choi and Miklós Csörgő. Asymptotic properties for increments of l^{∞}-valued Gaussian random fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):313–333, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Curry:2008:LPF

[438] Eva Curry. Low-pass filters and scaling functions for multivariable wavelets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):334–347, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Santos:2008:MFA

[439] F. Guillén Santos, V. Navarro, P. Pascual, and Agustí Roig. Monoidal functors, acyclic models and chain operads. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):348–378, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jorgensen:2008:FCM

[440] Peter Jørgensen. Finite Cohen– Macaulay type and smooth noncommutative schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):379–390, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Migliore:2008:GWL

[441] Juan C. Migliore. The geometry of the weak Lefschetz property and level sets of points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):391–411, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nguyen-Chu:2008:QCT

[442] G.-V. Nguyen-Chu. Quelques calculs de traces compactes et leurs transformées de Satake. (French) [Some calculations of compact traces and their Satake transforms]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):412–442, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shen:2008:CPF

[443] Z. Shen and G. Civi Yildirim. On a class of projectively flat metrics with constant flag curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):443–456, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Teplyaev:2008:HCF

[444] Alexander Teplyaev. Harmonic coordinates on fractals with finitely ramified cell structure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):457–480, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breuer:2008:HPR

[445] Florian Breuer and Bo-Hae Im. Heegner points and the rank of elliptic curves over large extensions of global fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):481–490, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bugeaud:2008:MFA

[446] Yann Bugeaud, Maurice Mignotte, and Samir Siksek. A multi-Frey approach to some multi-parameter families of Diophantine equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):491–519, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2008:MWN

[447] Chang-Pao Chen, Hao-Wei Huang, and Chun-Yen Shen. Matrices whose norms are determined by their actions on decreasing sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):520–531, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clark:2008:LBT

[448] Pete L. Clark and Xavier Xarles. Local bounds for torsion points on Abelian varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):532–555, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Draisma:2008:PSI

[449] Jan Draisma, Gregor Kemper, and David Wehlau. Polarization of separating invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):556–571, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hitrik:2008:NSP

[450] Michael Hitrik and Johannes Sjöstrand.

Non-selfadjoint perturbations of selfadjoint operators in two dimensions IIIa. one branching point. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):572–657, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mihailescu:2008:IPE

[451] Eugen Mihailescu and Mariusz Urbański.

Inverse pressure estimates and the independence of stable dimension for noninvertible maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):658–684, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Savu:2008:CEF

[452] Anamaria Savu. Closed and exact functions in the context of Ginzburg– Landau models. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):685–702, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toms:2008:SAA

[453] Andrew S. Toms and Wilhelm Winter.

Z-stable ASH algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):703–733, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baba:2008:GCQ

[454] Srinath Baba and Håkan Granath. Genus 2 curves with quaternionic multiplication. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):734–757, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bercovici:2008:HSP

[455] H. Bercovici, C. Foias, and C. Pearcy. On the hyperinvariant subspace problem. IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):758–789, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blasco:2008:TPC

[456] Laure Blasco. Types, paquets et changement de base: l’exemple de U(2, 1)(F_{0}). I. types simples maximaux et paquets singletons. (French) [Types, packages and base change: the case of U(2, 1)(F_{0}). I. Simple maximal types and singleton packets]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):790–821, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuwae:2008:MPS

[457] Kazuhiro Kuwae. Maximum principles for subharmonic functions via local semi-Dirichlet forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):822–874, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mare:2008:CQC

[458] Augustin-Liviu Mare. A characterization of the quantum cohomology ring of G/B and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):875–891, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neeb:2008:SCC

[459] Karl-Hermann Neeb and Friedrich Wagemann. The second cohomology of current algebras of general Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):892–922, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Okoh:2008:EKM

[460] F. Okoh and F. Zorzitto. Endomorphisms of Kronecker modules regulated by quadratic algebra extensions of a function field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):923–957, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2008:NCS

[461] Yichao Chen. A note on a conjecture of S. Stahl. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):958–959, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stahl:2008:EZS

[462] Saul Stahl. Erratum: “On the Zeros of Some Genus Polynomials”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):960, ????

2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [2].

Abrescia:2008:ADC

[463] Silvia Abrescia. About the defectivity of certain Segre–Veronese varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 961–974, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boca:2008:AAA

[464] Florin P. Boca. An AF algebra associated with the Farey tessellation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 975–1000, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deCornulier:2008:IGA

[465] Yves de Cornulier, Romain Tessera, and Alain Valette. Isometric group actions on Hilbert spaces: Structure of orbits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):1001–1009, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gale:2008:FCM

[466] José E. Galé and Pedro J. Miana. H^{∞} functional calculus and Mikhlin-type multiplier conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1010–1027, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hamblen:2008:LDG

[467] Spencer Hamblen. Lifting n-dimensional Galois representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1028– 1049, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huang:2008:APM

[468] Wen ling Huang and Peter Šemrl. Adjacency preserving maps on Hermitian matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1050–1066, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kariyama:2008:TUA

[469] Kazutoshi Kariyama. On types for unramified p-adic unitary groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 1067–1107, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lopez-Abad:2008:CTT

[470] J. Lopez-Abad and A. Manoussakis. A classification of Tsirelson type spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):1108–1167, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:2008:STB

[471] Michael Taylor. Short time behavior of solutions to linear and nonlinear Schrödinger equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):

1168–1200, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bahuaud:2008:HCS

[472] Eric Bahuaud and Tracey Marsh. Hölder compactification for some manifolds with pinched negative curvature near infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1201–1218, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baracco:2008:CEM

[473] Luca Baracco and Giuseppe Zampieri. CR extension from manifolds of higher type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1219–1239, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beliakova:2008:CCJ

[474] Anna Beliakova and Stephan Wehrli. Categorification of the colored Jones polynomial and Rasmussen invariant of links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1240–1266, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blake:2008:NRE

[475] Ian F. Blake, V. Kumar Murty, and Guangwu Xu. Nonadjacent Radix-τ expansions of integers in Euclidean imaginary quadratic number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 1267–1282, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ho:2008:RLP

[476] Kwok-Pun Ho. Remarks on Littlewood– Paley analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1283–1305, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muic:2008:TLT

[477] Goran Muić. Theta lifts of tempered representations for dual pairs (Sp_{2n}, O(V)). Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1306–1335, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olver:2008:MFL

[478] Peter J. Olver and Juha Pohjanpelto.

Moving frames for Lie Pseudo–Groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60 (??):1336–1386, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Romo:2008:DSS

[479] Fernando Pablos Romo. On ndimensional Steinberg symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??): 1387–1405, ???? 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ricotta:2008:HAP

[480] Guillaume Ricotta and Thomas Vidick.

Hauteur asymptotique des points de Heegner. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(??):1406–1436, ???? 2008. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2008:AII

[481] Anonymous. Author index — index des auteurs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 60(6):1437– 1440, December 2008. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behrend:2009:CCM

[482] Kai Behrend and Ajneet Dhillon. Connected components of moduli stacks of torsors via Tamagawa numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 3–28, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casanellas:2009:MRC

[483] M. Casanellas. The minimal resolution conjecture for points on the cubic surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):29–49, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2009:COB

[484] Huaihui Chen and Paul Gauthier. Composition operators on μ-bloch spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):50–75, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Christensen:2009:APA

[485] Lars Winther Christensen and Henrik Holm. Ascent properties of Auslander categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):76–108, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coskun:2009:ACK

[486] Izzet Coskun, Joe Harris, and Jason Starr. The ample cone of the Kontsevich moduli space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):109–123, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dijkstra:2009:CCE

[487] Jan J. Dijkstra and Jan van Mill. Characterizing complete Erdős space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 124–140, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:2009:LPM

[488] Ben Green and Sergei Konyagin. On the Littlewood problem modulo a prime. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):141–164, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laurent:2009:EDA

[489] Michel Laurent. Exponents of Diophantine approximation in dimension two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):165–189, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lu:2009:BHP

[490] Yufeng Lu and Shuxia Shang. Bounded Hankel products on the Bergman space of the polydisk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):190–204, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:2009:RNN

[491] M. Marshall. Representations of nonnegative polynomials, degree bounds and applications to optimization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 205–221, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nien:2009:KMG

[492] Chufeng Nien. Klyachko models for general linear groups of rank 5 over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):222–240, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Azamov:2009:OIS

[493] N. A. Azamov, A. L. Carey, P. G. Dodds, and F. A. Sukochev. Operator integrals, spectral shift, and spectral flow. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 241–263, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:2009:MAI

[494] J. P. Bell and K. G. Hare. On Z-modules of algebraic integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):264–281, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [550].

Bouya:2009:CIS

[495] Brahim Bouya. Closed ideals in some algebras of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 282–298, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dawson:2009:CSH

[496] Robert J. MacG. Dawson and Maria Moszyńska. Čebyšev sets in hyperspaces over R^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):299–314, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Enochs:2009:IRI

[497] E. Enochs, S. Estrada, and J. R. García Rozas. Injective representations of infinite quivers. applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 315–335, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garaev:2009:LSI

[498] M. Z. Garaev. The large sieve inequality for the exponential sequence λ^{[O(n^{15/14+o(1)})]} modulo primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 336–350, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:2009:MPH

[499] William Graham and Markus Hunziker.

Multiplication of polynomials on Hermitian symmetric spaces and Littlewood– Richardson coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):351–372, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKee:2009:IOW

[500] Mark McKee. An infinite order Whittaker function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):373–381, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miao:2009:UED

[501] Tianxuan Miao. Unit elements in the double dual of a subalgebra of the Fourier algebra A(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):382–394, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moriyama:2009:FAT

[502] Tomonori Moriyama. L-functions for GSp(2) × GL(2): Archimedean theory and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):395–426, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tadic:2009:RUC

[503] Marko Tadić. On reducibility and unitarizability for classical p-adic groups, some general results. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):427–450, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Valeriote:2009:SIP

[504] Matthew A. Valeriote. A subalgebra intersection property for congruence distributive varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):451–464, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woodford:2009:PPP

[505] Roger Woodford. On partitions into powers of primes and their difference functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):465–480, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banks:2009:UDF

[506] William D. Banks, Moubariz Z. Garaev, Florian Luca, and Igor E. Shparlinski. Uniform distribution of fractional parts related to pseudoprimes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 481–502, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baranov:2009:SBS

[507] Anton Baranov and Harald Woracek.

Subspaces of de Branges spaces generated by majorants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):503–517, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Belliard:2009:GUM

[508] Jean-Robert Belliard. Global units modulo circular units: Descent without Iwasawa’s main conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):518–533, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2009:GTN

[509] Chuan-Zhong Chen and Wei Sun. Girsanov transformations for nonsymmetric diffusions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):534–547, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Girouard:2009:FTC

[510] Alexandre Girouard. Fundamental tone, concentration of density, and conformal degeneration on surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 548–565, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:2009:CSC

[511] Ian Graham, Hidetaka Hamada, Gabriela Kohr, and John A. Pfaltzgraff. Convex subordination chains in several complex variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):566–582, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hajir:2009:APF

[512] Farshid Hajir. Algebraic properties of a family of generalized Laguerre polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):583–603, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hart:2009:FCC

[513] Joan E. Hart and Kenneth Kunen. First countable continua and proper forcing. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):604–616, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:2009:SIR

[514] Wook Kim. Square integrable representations and the standard module conjecture for general spin groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 617–640, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maeda:2009:CPI

[515] Sadahiro Maeda and Seiichi Udagawa.

Characterization of parallel isometric immersions of space forms into space forms in the class of isotropic immersions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):641–655, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCutcheon:2009:GPM

[516] Randall McCutcheon and Anthony Quas. Generalized polynomials and mild mixing. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):656–673, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pollack:2009:CRA

[517] David Pollack and Robert Pollack.

A construction of rigid analytic cohomology classes for congruence subgroups of SL_{3}(Z). Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):674–690, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yu:2009:PQS

[518] Xiaoxiang Yu. Prehomogeneity on quasi-split classical groups and poles of intertwining operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):691–707, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zelenyuk:2009:RHF

[519] Yevhen Zelenyuk. Regular homeomorphisms of finite order on countable spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):708–720, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calin:2009:SGS

[520] Ovidiu Calin, Der-Chen Chang, and Irina Markina. Subriemannian geometry on the sphere S^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):721–739, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caprace:2009:GFC

[521] Pierre-Emmanuel Caprace and Frédéric Haglund. On geometric flats in the CAT(0) realization of Coxeter groups and Tits buildings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):740–761, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DCruz:2009:HCF

[522] Clare D’Cruz and Tony J. Puthenpurakal. The Hilbert coefficients of the fiber cone and the a-invariant of the associated graded ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):762–778, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grbac:2009:RSS

[523] Neven Grbac. Residual spectra of split classical groups and their inner forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):779–806, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hong:2009:MOA

[524] Sunggeum Hong, Joonil Kim, and Chan Woo Yang. Maximal operators associated with vector polynomials of lacunary coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):807–827, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howard:2009:TGZ

[525] Benjamin Howard. Twisted Gross– Zagier theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):828–887, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Novik:2009:FRM

[526] Isabella Novik and Ed Swartz. Face ring multiplicity via CM-connectivity sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):888–903, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saliola:2009:FSA

[527] Franco V. Saliola. The face semigroup algebra of a hyperplane arrangement. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):904–929, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sidman:2009:PCA

[528] Jessica Sidman and Seth Sullivant. Prolongations and computational algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):930–949, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tange:2009:IIF

[529] Rudolf Tange. Infinitesimal invariants in a function algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):950–960, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernon:2009:TIO

[530] Florent Bernon. Transfert des intégrales orbitales pour les algèbres de Lie classiques. (French) [Transfer of orbital integrals for classical Lie algebras]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 961–1049, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bertin:2009:ECY

[531] Marie-Amélie Bertin. Examples of Calabi–Yau 3-folds of P^{7} with ρ = 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):1050–1072, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Griffiths:2009:RHK

[532] Ross Griffiths and Mikaël Lescop. On the 2-rank of the Hilbert kernel of number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1073–1091, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Irving:2009:MTF

[533] John Irving. Minimal transitive factorizations of permutations into cycles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):1092–1117, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pontreau:2009:PPS

[534] Corentin Pontreau. Petits points d’une surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1118–1150, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruan:2009:CMP

[535] Huo-Jun Ruan and Robert S. Strichartz.

Covering maps and periodic functions on higher dimensional Sierpinski gaskets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1151–1181, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strichartz:2009:PAP

[536] Robert S. Strichartz. Periodic and almost periodic functions on infinite Sierpinski gaskets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1182–1200, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arvanitoyeorgos:2009:IEM

[537] Andreas Arvanitoyeorgos, V. V. Dzhepko, and Yu. G. Nikonorov. Invariant Einstein metrics on some homogeneous spaces of classical Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1201–1213, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cilleruelo:2009:CLP

[538] Javier Cilleruelo and Andrew Granville. Close lattice points on circles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1214–1238, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davidson:2009:PRG

[539] Kenneth R. Davidson and Dilian Yang. Periodicity in rank 2 graph algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61 (??):1239–1261, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dong:2009:LLP

[540] Z. Dong. On the local lifting properties of operator spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1262–1278, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:2009:TBS

[541] Christopher Hoffman, Alexander E. Holroyd, and Yuval Peres. Tail bounds for the stable marriage of Poisson and Lebesgue. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1279–1324, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nien:2009:USM

[542] Chufeng Nien. Uniqueness of Shalika models. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1325–1340, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rivoal:2009:SPA

[543] Tanguy Rivoal. Simultaneous polynomial approximations of the Lerch function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1341–1356, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shen:2009:CLM

[544] Zhongmin Shen. On a class of Landsberg metrics in Finsler geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1357–1374, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spallone:2009:SDS

[545] Steven Spallone. Stable discrete series characters at singular elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1375–1382, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wambach:2009:IR

[546] Eric Wambach. Integral representation for U_{3} × GL_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):1383–1406, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Will:2009:TCR

[547] Pierre Will. Traces, cross-ratios and 2generator subgroups of SU(2, 1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1407–1436, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:2009:AII

[548] Anonymous. Author index — index des auteurs — for 2009 — pour 2009. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??): 1437–1440, ???? 2009. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http://cms.

                                        math.ca/cjm/v61/p1437.

Chiang:2010:EVD

[549] Yik-Man Chiang and Mourad E. H. Ismail. Erratum to: On value distribution theory of second order periodic ODEs, special functions and orthogonal polynomials [Cno. 2245272]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):261, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [348].

Bell:2012:CMA

[550] Jason P. Bell and Kevin G. Hare. Corrigendum to “On Z-modules of Algebraic Integers”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):254–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See